### v1

# Advancing Spatio-Temporal Processing in Spiking Neural Networks through Adaptation

Abstract
Efficient implementations of spiking neural networks on neuromorphic hardware promise orders of magnitude less power consumption than their non-spiking counterparts. The standard neuron model for spike-based computation on such neuromorphic systems has long been the leaky integrate-and-fire (LIF) neuron. As a promising advancement, a computationally light augmentation of the LIF neuron model with an adaptation mechanism experienced a recent upswing in popularity, caused by demonstrations of its superior performance on spatio-temporal processing tasks. The root of the superiority of these so-called adaptive LIF neurons however, is not well understood. In this article, we thoroughly analyze the dynamical, computational, and learning properties of adaptive LIF neurons and networks thereof. We find that the frequently observed stability problems during training of such networks can be overcome by applying an alternative discretization method that results in provably better stability properties than the commonly used Euler-Forward method. With this discretization, we achieved a new state-of-the-art performance on common event-based benchmark datasets. We also show that the superiority of networks of adaptive LIF neurons extends to the prediction and generation of complex time series. Our further analysis of the computational properties of networks of adaptive LIF neurons shows that they are particularly well suited to exploit the spatio-temporal structure of input sequences. Furthermore, these networks are surprisingly robust to shifts of the mean input strength and input spike rate, even when these shifts were not observed during training. As a consequence, high-performance networks can be obtained without any normalization techniques such as batch normalization or batch-normalization through time.
1 Introduction
Spiking neural networks (SNNs) have emerged as a viable biologically inspired alternative to artificial neural network (ANN) models [maass1997networks]. In contrast to ANNs, where neurons communicate analog numbers, neurons in SNNs communicate via digital pulses, so-called spikes. This event-based communication resembles the communication of neurons in the brain and enables highly energy efficient implementations in neuromorphic hardware [schemmel2010wafer, furber2014spinnaker, merolla2014million, davies2018loihi, davies2021advancing]. Recent advances in SNN research have shown that SNNs can be trained in a similar manner as ANNs using backpropagation through time (BPTT), leading to highly accurate models [leeTrainingDeepSpiking2016, slayer].
The dominant spiking neuron model used in SNNs is the leaky integrate and fire (LIF) neuron [gerstner2014neuronal]. The LIF neuron has a single state variable $u(t)$ that represents the membrane potential of a biological neuron. Incoming synaptic currents are integrated over time in a leaky manner on the time scale of tens of milliseconds. Once the membrane potential reaches a threshold $\vartheta$, the membrane potential is reset and the neuron emits a spike (i.e., its output is set to $1$). The leaky integration property of the LIF neuron model reproduces the sub-threshold behavior of so-called excitability class 1 neurons in the brain (Fig. 1a, left) [hodgkinLocalElectricChanges1948, izhikevichNeuralExcitabilitySpiking2000a].
Figure 1: The adLIF neuron with membrane potential oscillation and spike-frequency adaptation. a) Neurons in the brain have been classified into two excitability classes, integrators (class 1) and resonators (class 2). While resonators show membrane potential oscillations, integrators do not. b) Membrane potential oscillation in response to an input pulse. The period $P$ and decay $r$ are sufficient to fully characterize the oscillating spike response. c) Example of a neuron without (top) and with (bottom) spike frequency adaptation (SFA). Dotted arrow illustrates the feed-back of output spikes to the neuron membrane state in the adLIF model. d) Adaptive LIF (adLIF) neurons (see Eq. (2)) differ in 2 features from vanilla LIF neurons: membrane potential oscillations and SFA. e) Impulse response functions for different parameterizations of adLIF neurons. For $a=0$ (green), no oscillations occur and the spike response reduces to leaky integration.
A second class of neurons called excitability class 2 neurons (Fig. 1a, right) exhibit more complex dynamics with sub-threshold membrane potential oscillations (Fig. 1b) and spike frequency adaptation (SFA, Fig. 1c). Such complex dynamics cannot be modelled with the single state variable $u(t)$ of the LIF neuron. Pioneering work has shown that these behaviors can be reproduced by a simple extension of the LIF neuron model that adds a second state variable to the neuron dynamics which interacts with the membrane potential in a — typically negative — feedback loop [izhikevich2001resonate]. Neuron models of this type are called adaptive LIF neurons.
With the growing interest in SNNs for neuromorphic systems, researchers have started to train recurrent SNNs (RSNNs) consisting of adaptive LIF neurons with BPTT on spatio-temporal processing tasks. First results were based on neuron models that implement a threshold adaptation mechanism, where the second state variable is a dynamic threshold $\vartheta(t)$ [bellec2018long, salaj2021spike, gangulySpikeFrequencyAdaptation2024a]. Each spike of the neuron leads to an increase of this threshold, which implements the negative feedback loop mentioned above and leads to SFA (Fig. 1c). The performance of these models clearly surpassed those achieved with networks of LIF neurons while being highly efficient on neuromorphic hardware with orders of magnitudes energy savings when compared to implementations on CPUs or graphical processing units (GPUs) [davies2021advancing].
While threshold adaptation implements SFA, the resulting neuron model still performs a leaky integration of input currents and does not exhibit the typical sub-threshold membrane potential oscillations of class 2 neurons. Hence, more recent work considered networks of neurons with a form of adaptation often referred to as sub-threshold or current-based adaptation. The second state variable $w(t)$ is interpreted as negative adaptation current that is increased not only by neuron spikes but also by the sub-threshold membrane potential itself. This sub-threshold feedback leads to complex oscillatory membrane potential dynamics (Fig. 1b). Interestingly, simulation studies have shown that SNNs equipped with sub-threshold adaptation achieve significantly better performances than SNNs with threshold adaptation [bittar2022surrogate, gangulySpikeFrequencyAdaptation2024a, higuchi2024balanced].
Although the converging evidence suggests that networks of adaptive LIF neurons are superior to LIF networks for neuromorphic applications, there are still many questions open. First, to achieve top performance, usually all neuron parameters are trained together with the synaptic weights. Changes of the neuron parameters however can quickly lead to unstable models which disrupts training. To avoid instabilities, parameter bounds have to be defined and fine-tuned. If these bounds are too wide, the network can become unstable, if they are too narrow, one cannot utilize the full computational expressivity. We show that this problem is not inherent to the neuron model but rather caused by the standard discrete-time formulation of the continuous neuron dynamics which is based on the Euler-Forward discretization method. Our thorough theoretical analysis reveals that an alternative discretization method, the so-called Symplectic-Euler discretization, provably leads to a much more stable discrete adaptive neuron model, where the full expressivity of adaptive neurons can be utilized at the same computational costs. Using this insight, we demonstrate the power of adaptation by showing that our improved adaptive RSNNs outperform the state-of-the-art on spiking speech recognition datasets as well as an ECG dataset. We then show that the superiority of adaptive RSNNs is not limited to classification tasks but extends to the prediction and generation of complex time series.
Second, there is a lack of understanding why sub-threshold adaptation is so powerful in RSNNs. We thoroughly analyze the computational dynamics in single adaptive LIF neurons, as well as in networks of such neurons. Our analysis suggests that adaptive LIF neurons are especially capable of detecting temporal changes in input spike density, while being robust to shifts of total spike rate. Hence, adaptive RSNNs are well suited to analyze the temporal properties of input sequences.
Third, high-performance SNNs are usually trained using normalization techniques such as batch normalization or batch normalization through time [batchnorm, kim2021revisitingbatchnormalizationtraining, deckers2024co, vicente-solaSpikingNeuralNetworks2023]. These methods however complicate the training process and the implementation of networks on neuromorphic hardware. We show that adaptation has a previously unrecognized benefit on network optimization. Since adaptation inherently stabilizes network activity, we hypothesized that explicit normalization is not necessary in adaptive RSNNs. In fact, all our results were obtained without explicit normalization techniques. We test this hypothesis and show that in contrast to LIF networks, networks of adaptive LIF neurons can tolerate substantial shifts in the mean input strength as well as substantial levels of background noise even when these perturbations were not observed during training.
2 Results
2.1 Adaptive LIF neurons
The leaky integrate-and-fire (LIF) neuron model [gerstner2014neuronal] evolved as the gold standard for spiking neural networks due to its simplicity and suitability for low-power neuromorphic implementation [rathi2023exploring]. The continuous-time equation for the membrane potential of the LIF neuron at time $t$ is given by
$\tau_{u}\dot{u}=-u(t)+I(t),$ (1)
where $\tau_{u}$ is the membrane time constant, and $I(t)=\sum_{i}\theta_{i}x_{i}(t)$ the input current composed of the sum of neuron inputs $x_{i}(t)$ scaled by the corresponding synaptic weights $\theta_{i}$. A dot above a variable denotes its derivative with respect to time. If the membrane potential $u(t)$ crosses the spike threshold $\vartheta$ from below, a spike is emitted and $u$ is reset to the reset potential. In the absence of input, $u(t)$ decays exponentially to zero. The LIF neuron equation models so-called integrating class 1 neurons in the brain (Fig. 1a), which are integrating incoming currents in a leaky manner. The simple first-order dynamics however does not allow the LIF model to account for another class of neurons frequently occurring in the brain: resonating/oscillating class 2 neurons (Fig. 1a,b). In contrast to integrators, such neurons exhibit oscillatory behavior in response to stimulation, giving rise to interesting properties entirely neglected by LIF neurons. Such oscillatory behavior is often modelled by adding a second time-varying variable — the adaptation current $w(t)$ — to the neuron state [izhikevich2001resonate, brunelFiringrateResonanceGeneralized2003, bittar2022surrogate, deckers2024co]. The resulting neuron model, which we refer to as adaptive leaky integrate-and-fire (adLIF) neuron, has significant advantages over LIF neurons in terms of feature detection capabilities and gradient propagation properties, as we show in the next few sections. The adLIF model is described in terms of two coupled differential equations
$\tau_{u}\dot{u}$ $=-u(t)+I(t)-w(t)$ (2)
$\tau_{w}\dot{w}$ $=-w(t)+au(t)+bz(t),$ (3)
where $\tau_{w}$ is the adaptation time constant and $a$ and $b$ are adaptation parameters, defining the behavior of the neuron. When comparing the LIF equation (1) with equation (2), we see that the latter resembles the LIF dynamics where the adaptation current $w(t)$ is subtracted, with its dynamics defined in equation (3).
The parameter $a\in\mathbb{R}$ scales the coupling of the membrane potential $u(t)$ with the adaptation current $w(t)$. The negative feedback loop between $u(t)$ and $w(t)$ defined by Equations (2) and (3) leads to oscillations of the membrane potential for large enough $a$, see Fig. 1b. The oscillation can be characterized by the decay rate $r$ and the period $P=\frac{1}{f}$ given by the inverse of the intrinsic frequency $f$. As we discuss later in the manuscript, $f$ characterizes the frequency tuning of the neuron, whereas the decay rate $r$ is an indicator of its stability and time scale.
Figure 2: The adLIF neuron model shows frequency-selective resonating behavior. a) Voltage response (root mean squared membrane potential over $10$ seconds) of a LIF neuron in response to tonic spike input of different rates. b) Membrane potential response of an adLIF neuron with intrinsic frequency of $f\approx 60$ Hz to an input spike triplet at $60$ Hz (left) and $100$ Hz (right). c) Voltage response of the same adLIF neuron as in panels b and c for tonic spiking input at different rates. d) A $10$ Hz sinusoidal input signal (top) is encoded as an input spike train (middle) through spike frequency modulation (SFM), see main text for details. Membrane potential response of an adLIF neuron with intrinsic frequency $f\approx 10$ Hz (bottom). e) Same as panel e, but for a sinusoidal input at $7$ Hz. f) Voltage response of an adLIF neuron to SFM-encoded sinusoidal input at various frequencies. g) Same as panel f, but for a LIF neuron. See Methods for parameters and input generation.
The parameter $b\geq 0$ weights the feed-back from the neuron’s output spike $z(t)$ onto the adaptation variable $w(t)$. Hence, each spike has an inhibitory effect on the membrane potential, which leads to spike frequency adaptation (SFA) (gangulySpikeFrequencyAdaptation2024a), see Fig. 1c. We refer to this auto-feed-back governed by parameter $b$ as spike-triggered adaptation in the following. SFA has also been implemented directly using an adaptive firing threshold that is increased with every output spike yinAccurateEfficientTimedomain2021; lsnn_igi. In contrast to the adLIF model, these models do not exhibit membrane potential oscillations.
The adLIF model combines both membrane potential oscillations and SFA in one single neuron model (Fig. 1d). Depending on the parameters, adLIF neurons can exhibit oscillations of diverse frequencies and decay rates (Fig. 1e), and are equivalent to LIF neurons for $a,b=0$, where neither oscillations nor spike-triggered adaptation occur. A reduced variant of the adLIF neuron is given by the resonate-and-fire neuron izhikevich2001resonate.
Originally developed to efficiently replicate firing patterns of biological neurons, the adLIF model recently gained attention due to significant performance gains over vanilla LIF neurons in several benchmark tasks, despite its little computational overhead (bittar2022surrogate; gangulySpikeFrequencyAdaptation2024a; deckers2024co). In particular, gradient-based training of networks of adLIF neurons on spatio-temporal processing tasks appears to synergize well with oscillatory dynamics. However, these empirical findings are so far not accompanied by a good understanding of the reasons for this superiority.
When comparing the responses of the LIF and adLIF neuron, an important computational consequence of membrane potential oscillations has been noted: In contrast to the LIF neuron, which responds with higher amplitude of $u$ to higher input spike frequency (Fig. 2a), the adLIF neuron is most strongly excited if the frequency of input spikes matches the intrinsic frequency $f$ of the neuron (Fig. 2b-c, see also izhikevich2001resonate; brunelFiringrateResonanceGeneralized2003; higuchi2024balanced). To demonstrate this resonance phenomenon, we show the membrane potential of an adLIF neuron with intrinsic frequency $f=60$ Hz for an input spike triplet exactly at this intrinsic frequency $f$ (Fig. 2b, left), compared to a spike triplet of higher rate (Fig. 2b, right). The resulting amplitude of the membrane potential $u$ is higher in the former case, indicating resonance. Fig. 2c shows that the neuron exhibits a frequency selectivity specifically for its intrinsic frequency $f$.
We took this analysis a step further and asked whether this resonance could account for frequencies in the input spike train that are not directly encoded by spike rate, but rather by slow changes of the spike rate over time, a coding scheme previously termed spike frequency modulation (SFM) sfmwang. As a guiding example, we encoded a slow-varying sinusoidal signal as a spike train, where the magnitude of the signal at a certain time is given by the local spike rate, shown in Fig. 2d. The spike rate thereby varied between 00 Hz and $200$ Hz, whereas the underlying, encoded sinus signal oscillated with a constant frequency of $10$ Hz. Again, we see increased response of the membrane potential $u$ over time in the case of the $10$ Hz input compared to a slower $7$ Hz sinus signal (Fig. 2d,e), due to resonance with the adLIF neuron, see also Fig. 2f. In contrast, the corresponding membrane voltage response amplitude of a LIF neuron is almost indifferent to the intrinsic frequency of the underlying sinus input, see Fig. 2g. This shows that in contrast to the LIF neuron, the adLIF neuron model is sensitive to the longer-term temporal structure, i.e. variation of the input signal. In Section Computational properties of adLIF networks, we highlight the importance of this frequency-dependence of neuron responses as a key ingredient for the powerful feature detection capabilities of networks of adLIF neurons.
2.2 The Symplectic-Euler discretized adLIF neuron
In the previous section, we defined the LIF and adLIF neuron models via continuous-time, ordinary differential equations. In practice, it is however standard to discretize the continuous-time dynamics of the spiking neuron model. This allows not only to use powerful auto-differentiation capabilities of machine learning software packages such as TensorFlow (tensorflow2015-whitepaper) or PyTorch(anselPyTorchFasterMachine2024), but also for implementation of such neuron models in discrete-time operating neuromorphic hardware orchardEfficientNeuromorphicSignal2021. Discretization of the LIF neuron model Eq. (1) is straight-forward. In contrast, for the adLIF model, the interdependency of the two state variables during a discrete time step $\Delta t$ cannot be taken into account exactly in a simple manner (the exact solution involves a matrix exponential). Nevertheless, for efficient simulation and hardware implementations, simple update equations are needed. Therefore, approximate discrete update equations for the membrane potential $u$ and the adaptation current $w$ are usually obtained by the Euler-Forward method bittar2022surrogate. In the following we analyze discretization methods for adLIF neurons through the lens of dynamical systems analysis. We find that the Euler-Forward method is problematic, and propose the utilization of a more stable alternative discretization method.
A common approach to study dynamical systems is through the state-space representation, which recently gained popularity in the field of deep learning guCombiningRecurrentConvolutional2021; guMambaLinearTimeSequence2023a. Re-formulation of spiking neuron models in a canonical state-space representation provides a convenient unified way to study their dynamical properties. The continuous-time equations of the adLIF neuron Eq. (2), (3) can be re-written in such a state-space representation as a 2-dimensional linear time-invariant (LTI) system with state vector $\bm{s}$ as
$\dot{\bm{s}}(t)=\begin{pmatrix}\dot{u}(t)\\ \dot{w}(t)\end{pmatrix}$ $=A\bm{s}(t)+B\bm{x}(t)$
$=\underbrace{\begin{pmatrix}-\frac{1}{\tau_{u}}&-\frac{1}{\tau_{u}}\\ \frac{a}{\tau_{w}}&-\frac{1}{\tau_{w}}\end{pmatrix}}_{A}\begin{pmatrix}u(t)\\ w(t)\end{pmatrix}+\underbrace{\begin{pmatrix}\frac{1}{\tau_{u}}&0\\ 0&\frac{b}{\tau_{w}}\\ \end{pmatrix}}_{B}\begin{pmatrix}I(t)\\ z(t)\end{pmatrix},$ (4)
with system matrix $A$ and input matrix $B$. This equation only describes the sub-threshold dynamics of the neuron (i.e., it holds as long as the threshold is not reached). The reset can be accounted for by the threshold condition: When the voltage crosses the firing threshold $\vartheta$, an output spike is elicited and the neuron is reset. The goal of discretization is to obtain discrete-time update equations of the form
$\bm{s}[k]=\begin{pmatrix}u[k]\\ w[k]\end{pmatrix}=\bar{A}\bm{s}[k-1]+\bar{B}\bm{x}[k],$ (5)
where $f[k]$ denotes the value of state variable $f$ at discrete time step $k$, i.e., $f[k]\equiv f(\Delta tk)$ for discrete time increment $\Delta t$ and integer-valued $k>0$. Here, $\bar{A}$ and $\bar{B}$ denote the state and input matrix of the discrete time system respectively. In the SNN literature, the most commonly used approach to obtain the discrete approximation to the continuous system from Eq. (4) is the Euler-Forward method bittar2022surrogate; deckers2024co; higuchi2024balanced, which results in update equations
$\hat{u}[k]$ $=\alpha u[k-1]+(1-\alpha)\left(-w[k-1]+I[k]\right)$ (6a)
$w[k]$ $=\beta w[k-1]+(1-\beta)(au[k-1]+bS[k]),$ (6b)
where $\hat{u}$ denotes the membrane potential before the reset is applied, $\alpha=1-\frac{\Delta t}{\tau_{u}}$, and $\beta=1-\frac{\Delta t}{\tau_{w}}$. The spike output of the neuron is given by
$S[k]=\begin{cases}1&\text{if }\hat{u}[k]>u_{\mathrm{th}}\\ 0&\text{otherwise.}\end{cases}$ (7)
Finally, $u[k]$ is obtained by applying the reset to $\hat{u}$ via
$u[k]=\hat{u}[k]\cdot(1-S[k]).$ (8)
This Euler-Forward discretization yields the discrete state-space matrices
$\bar{A}_{\mathrm{EF}}=\begin{pmatrix}\alpha&-(1-\alpha)\\ a(1-\beta)&\beta\end{pmatrix}\hskip 30.0pt\bar{B}_{\mathrm{EF}}=\begin{pmatrix}(1-\alpha)&0\\ 0&b(1-\beta)\\ \end{pmatrix}$ (9)
for $\bar{A}$ and $\bar{B}$ in Eq. (5). In practice (see for example bittar2022surrogate), the coefficients $\alpha$ and $\beta$ are often replaced by exponential decay terms $\alpha=\exp\left(-\frac{\Delta t}{\tau_{u}}\right)$ and $\beta=\exp\left(-\frac{\Delta t}{\tau_{w}}\right)$ akin to the LIF discretization, as it is exact in the latter case. However, for adLIF neurons, the Euler-Forward approximation is quite imprecise which can quickly result in unstable and diverging behavior of the system, as we will show below. A better approximation is given by the bilinear discretization method, a standard method also used in state space models guCombiningRecurrentConvolutional2021, which is however computationally more demanding. An alternative is the Symplectic-Euler (SE) method geometric_numerical_integration that has previously been used in non-spiking oscillatory systems effenbergerFunctionalRoleOscillatory2023b. We found that the Symplectic-Euler (SE) discretization provides major benefits in terms of stability, expressivity, and trainability of the adLIF neuron, while being computationally as efficient as Euler-Forward. The SE method has been shown to preserve the energy in Hamiltonian systems, a desirable property of a discretization of such systems geometric_numerical_integration. As we show below, the improved stability of the SE method still applies to the adLIF neuron model, even though it is non-Hamiltonian. The SE method is similar to the Euler-Forward method, with the only difference that one computes the state variable $w[k]$ from $u[k]$ instead of $u[k-1]$, resulting in the discrete dynamics
$\hat{u}[k]$ $=\alpha u[k-1]+(1-\alpha)\left(-w[k-1]+I[k]\right)$ (10a)
$w[k]$ $=\beta w[k-1]+(1-\beta)(au[k]+bS[k]).$ (10b)
We refer to this neuron model as the SE-adLIF model in order to distinguish it from the Euler-Forward discretized model. Note that the reset mechanism from Eq. (8) is applied to obtain $u[k]$ from $\hat{u}[k]$ before computing $w[k]$. While it is also possible to apply the reset after computing $w[k]$, we found that the above described way yields the best performance. For the sub-threshold dynamics, this leads to update matrices (see Section 4.2 in Methods)
$\bar{A}_{\mathrm{SE}}=\begin{pmatrix}\alpha&-(1-\alpha)\\ a(1-\beta)\alpha&\quad\beta-a(1-\beta)(1-\alpha)\\ \end{pmatrix}\hskip 30.0pt\bar{B}_{\mathrm{SE}}=\begin{pmatrix}(1-\alpha)&0\\ 0&b(1-\beta)\\ \end{pmatrix}$ (11)
for the discrete state-space formulation given by Eq. (5).
2.3 Stability analysis of discretized adLIF models
Training of networks of adLIF neurons often suffers from stability issues, in particular when the neuron parameters such as the time constants $\tau_{w}$, $\tau_{u}$ or the adaptation parameters $a$, $b$ are trained. The decay rate $r$ (see Fig. 1b) determines the asymptotic stability of the neuron dynamics: A decay rate $r>1$ results in a neuron in which the state vector grows exponentially over time, leading to an unstable and diverging system, whereas a neuron with decay rate $r<1$ exhibits stable dynamics where the state vector exponentially decays to a resting state equilibrium. Since the continuous-time adLIF neuron model Eq. (2) is inherently stable for $a\geq-1$ (see Section 4.3 in Methods for proof), it is a desirable property of a discretization method to preserve this stability for all possible parameterizations. However, when the model is time-discretized, the discrete-time decay rate can deviate from its continuous counterpart due to discretization errors, leading to stability properties that deviate from the continuous-time system. This is shown empirically in Fig. 3a. We instantiated $1{,}000$ different neurons for both discretization methods, Euler-Forward and SE, as well as the continuous model, in a grid-like manner over a reasonable parameter range and plot their calculated frequencies and decay rates in Fig. 3a. While the continuous system is stable for all considered parameter combinations (middle panel), the Euler-Forward approximation (left panel) is unstable for many parameterizations (decay rate $r>1$). In contrast, for the SE discretization (right panel), all parameter combinations resulted in stable neuron dynamics ($r<1$). These empirical results show that the SE discretization more closely follows the stability properties of the continuous model, whereas the Euler-Forward method deviates drastically from both. How can this discrepancy be explained?
When analyzing the discretized neurons, one has to calculate $f$ and $r$ directly from the discrete system by calculating the eigenvalues $\lambda_{1,2}$ of the state transition matrix $\bar{A}$. This allows to study the behavior and the stability of the adLIF neuron model for different discretizations. Two cases have to be differentiated: If the eigenvalues are complex, the membrane potential exhibits oscillations, whereas if they are real, no oscillations occur and the neuron behaves similar to a LIF neuron. In the complex case, we can write the eigenvalues in polar form as $\lambda_{1,2}=re^{\pm j\phi}$, where $j$ denotes the imaginary unit. Hence, the eigenvalues are complex conjugates, the decay rate $r$ is given by their modulus (see Fig. 3b) and $\phi$ is obtained as the argument of $\lambda_{1}$ ($\text{arg}(\lambda_{2})=-\phi$). Intuitively, the angle $\phi$ is the rotation of the neuron state with each time step $\Delta t$ in radians, and hence determines the frequency of the oscillation. One thus obtains the intrinsic frequency $f$ in Hertz as $f=\frac{\phi}{2\pi\Delta t}$. In the case of real eigenvalues, $r$ is given by the magnitude of the largest eigenvalue. AdLIF neurons can thereby represent underdamped (complex eigenvalues), critically damped (equal real eigenvalues), and overdamped (non-equal real eigenvalues) systems via different parameterizations. Note, that only in the underdamped case the neuron can oscillate.
Figure 3: Stability of adLIF discretizations. a) Intrinsic frequency $f$ and per-timestep decay rate $r$ of $1{,}000$ different parameterizations of adLIF neurons for Euler-Forward discretization (left), SE (right), and the continuous model (middle). The horizontal dotted line at $r=1.0$ marks the stability bound. Instances above this line diverge due to exponential growth. Parameter ranges are uniformly distributed over the intervals $a\in[0,120]$, $\tau_{u}\in[5,25]$ ms and $\tau_{w}\in[60,300]$ ms. b) Eigenvalues of $\bar{A}_{\mathrm{EF}}$ (left) and $\bar{A}_{\mathrm{SE}}$ (right) plotted in the complex plane for fixed $\tau_{u}=25$ ms, $\tau_{w}=60$ ms and varying $a\in[10,800]$. Decay rate $r$ as modulus of the eigenvalue $\lambda_{1}$ and angle $\phi$ as argument of $\lambda_{1}$ are shown for $a=282$, marked with * and ** for EF and SE respectively. The grey half-circle denotes the stable region of $r\leq 1$. c) Relationship of parameter $a$ to intrinsic frequency $f$ (top) and decay rate $r$ (bottom) for the same $\tau_{u}$ and $\tau_{w}$ as in panel b. Points with * and ** denote the corresponding eigenvalues from panel b. Recall the linear relationship $f=\frac{\phi}{2\pi\Delta t}$ between angle $\phi$ and $f$ of the discrete models. Horizontal grey line in bottom panel denotes stability boundary of $r=1$. Values for $r$ of continuous model ($r=0.897$) and SE-adLIF ($r=0.887$) are constant w.r.t. $a$ (SE-adLIF values for $r$ not visible due to near-perfect fit to the continuous model). d) Maximum admissible frequency for stable dynamics for Euler-Forward discretization with $\Delta t=1$ ms over different values of $\tau_{u}$ and $\tau_{w}$, where $a$ is set to the maximum stable value $a_{\mathrm{max}}$ (see main text and Section 4.7 in Methods).
As described above, the eigenvalues $\lambda_{1,2}$ of state transition matrices $\bar{A}_{\mathrm{EF}}$ and $\bar{A}_{\mathrm{SE}}$ are determining the stability of the discrete neurons. We can directly observe the origin of instability for the EF-adLIF by plotting eigenvalues for different neuron parameters in the complex plane. In Fig. 3b, we show some eigenvalues of $\bar{A}_{\mathrm{EF}}$ and $\bar{A}_{\mathrm{SE}}$ in the case of fixed time constants $\tau_{u}$ and $\tau_{w}$ and varying parameter $a$. In the complex plane, the stability boundary appears as circle, separating the stable ($r<1$, grey area) from the unstable region ($r>1$). Our analysis in Section 4.4 in Methods shows that for the Euler-Forward method, for fixed time constants $\tau_{u}$ and $\tau_{w}$, the real part $\Re(\lambda_{1,2})$ of these eigenvalues is constant and strictly positive with respect to $a$, such that the eigenvalues are aligned along a vertical line in the right half-plane of the complex plane (see Fig. 3b, left panel). As $a$ increases, the imaginary part increases and so does the decay rate $r$. Hence, for the Euler-Forward discretized neuron, the eigenvalues overshoot the stability boundary $r=1$ for increasing $a$. This leads to a drastically reduced range of the angle $\phi$ and therefore a reduced range of the intrinsic frequency $f$ where the neuron is stable. In contrast, for the SE discretized model, the parameter $a$ controls only the angle $\phi$ of the eigenvalues (see Section 4.5 in Methods) and hence the intrinsic frequency $f$, but not the decay rate $r$. The decay rate is given by $r=\sqrt{\alpha\beta}$ (see Eq. (55)) and is hence guaranteed to stay within the stability bound $r<1$ for all $\tau_{w}>0$ and $\tau_{u}>0$, see Fig. 3b, right panel and Fig. 3c.
We analytically calculated stability bounds for both the Euler-Forward and SE discretization, see Methods Sections 4.4 and 4.5. For each given tuple of time constants $\tau_{u}$ and $\tau_{w}$ we calculated a corresponding $a_{\mathrm{max}}$, that is, the maximum value of the parameter $a$ for which the model is still stable. This analysis shows that the SE discretization allows the neuron to utilize the full frequency bandwidth up to the Nyquist frequency at $\frac{1}{2\Delta t}$, at which aliasing occurs. Since we used a discretization time step of $\Delta t=1$ ms for Fig. 3, the Nyquist frequency is $500$ Hz. Theorem 2.1 below summarizes the full frequency coverage of SE-adLIF and the stability within this frequency range (see Methods, Section 4.6 for a proof).
Theorem 2.1.
Let $(\tau_{u},\tau_{w},a)$ be the parameters of an SE-adLIF neuron according to Eq. (11). For any frequency $f\in(0,f_{N}]$ where $f_{N}=\frac{1}{2\Delta t}$ is the Nyquist frequency, and for any $\tau_{u},\tau_{w}>0$, there exists a unique parameter $a$ such that the neuron has intrinsic frequency $f$. For any such parameter combination, the neuron in the sub-threshold regime is asymptotically stable with decay rate $r=\sqrt{\alpha\beta}<1$ where $\alpha=e^{-\frac{\Delta t}{\tau_{u}}}$ and $\beta=e^{-\frac{\Delta t}{\tau_{w}}}$.
The upper frequency bound of adLIF neurons using the Euler-Forward discretization is illustrated in Fig. 3d. We can observe that for the Euler-Forward method, the maximum admissible frequency for stable dynamics converges toward zero as $\tau_{u}$ and $\tau_{w}$ increase (see Sections 4.7 in Methods for proof).
The difference in the stability properties has direct implications on network training: Firstly, since the neuron parameters $\tau_{u}$, $\tau_{w}$, $a$, and $b$ are usually trained, parameter ranges for Euler-Forward discretized adLIF neurons need to be constrained carefully to avoid stability issues. A similar issue has been observed in higuchi2024balanced and deckers2024co. Secondly, since SE discretization allows adLIF neurons to obtain oscillations of the full frequency bandwidth, they are theoretically capable of tuning to all possible frequencies present in the input.
2.4 Improved performance of SE-discretized adaptive RSNNs
In the previous sections, we discussed the theoretical advantages of SE discretization over the more commonly used Euler-Forward discretization of the adLIF neuron model. From these theoretical considerations however, it is not directly evident whether this method provides any advantages in a realistic task setup. We therefore evaluated how recurrent networks of adLIF neurons compete against classical vanilla LIF networks as well as networks of adaptive neurons that have previously been considered in the literature: a model with threshold adaptation (ALIF) yinAccurateEfficientTimedomain2021, a constrained variant of the adLIF neuron model, similar to the model in this study, but with differences in discretization and neuron formulation (cAdLIF) deckers2024co, another adLIF network but with batch normalization (RadLIF) bittar2022surrogate, a feed-forward model with delays implemented as temporal convolutions (DCLS-Delays) hammouamri2023learning, and the balanced resonate-and-fire neuron model (BHRF) higuchi2024balanced, a variant of the resonate-and-fire neuron izhikevich2001resonate where output spikes do not disrupt the phase of the membrane potential oscillation. To that end, we constructed a recurrently connected SNN composed of one or two layers (depending on the task) of adLIF neurons, followed by a layer of leaky integrator (LI) neurons to provide a real-valued network output. We trained the SNN using BPTT with surrogate gradients leeTrainingDeepSpiking2016; bellec2018long; neftci2019surrogate; slayer. We used a dropout rate of $15\%$ except noted otherwise, but otherwise no regularization, normalization, or data augmentation methods. The trained parameters included the synaptic weights $\theta$, as well as all neuron parameters $a$, $b$, $\tau_{u}$ and $\tau_{w}$, which were not shared across neurons such that each neuron could have individual parameter values. Neurons were initialized heterogeneously, such that for each neuron the initial values of these parameters were chosen randomly from a uniform distribution over a pre-defined range. Heterogeneity has previously been shown to improve the performance of SNNs perez-nievesNeuralHeterogeneityPromotes2021. We applied a reparametrization technique for the training of time constants $\tau_{u}$ and $\tau_{w}$ and parameters $a$ and $b$, see Methods for details.
We evaluated the networks on two commonly used audio benchmark datasets: Spiking Heidelberg Digits (SHD) cramerHeidelbergSpikingData2022 and Spiking Speech Commands (SSC) cramerHeidelbergSpikingData2022 , as well as an ECG dataset previously used to test ALIF neurons yinAccurateEfficientTimedomain2021. In Table 1 we report the test accuracy on the corresponding test sets. SHD does not define a dedicated validation set and previous work reported performances for networks validated on the test set, which is methodologically not clean. We therefore report results for two validation variants for SHD: with validation on the test set (to ensure comparability) and with validation on a fraction of the training set.
Model Rec. #Params #Runs Test Acc. [%]
SHD LIF $0.45$M $10$ $90.27\pm 0.73\%$
ALIF yinAccurateEfficientTimedomain2021 $0.14$M $1$ $90.4\%$
cAdLIF deckers2024co $38.7$k $10$ $94.19\%$
RadLIF bittar2022surrogate $3.9$M - $94.62\%$
DCLS-Delays hammouamri2023learning $0.2$M $10$ $95.07\pm 0.24\%$
EF-adLIF (2 layers) $0.45$M $20$ $94.68\pm 0.57\%$
SE-adLIF (1 layer) $37.5$k 20 $94.59\pm 0.27\%$
SE-adLIF (2 layers) $0.45$M 20 $95.81\pm 0.56\%$
SHD* BHRF higuchi2024balanced $0.1$M $5$ $92.7\pm 0.7\%$
SE-adLIF (1 layer) $37.5$k 20 $93.18\pm 0.74\%$
SE-adLIF (2 layers) $0.45$M 20 $93.79\pm 0.76\%$
SSC LIF $1.7$M $1$ $75.23\%$
ALIF yinAccurateEfficientTimedomain2021 $0.78$M $1$ $74.2\%$
RadLIF bittar2022surrogate $3.9$M - $77.4\%$
cAdLIF deckers2024co $0.35$M $1$ $77.5\%$
DCLS-Delays hammouamri2023learning $2.5$M $5$ $80.69\pm 0.21\%$
EF-adLIF (2 layers) $1.6$M $5$ $79.80\pm 0.18\%$
SE-adLIF (2 layers) $1.6$M $5$ $80.44\pm 0.26\%$
ECG LIF (1 layer) $1.8$k 5 $77.8\pm 1.85\%$
ALIF yinAccurateEfficientTimedomain2021 $1.8$k $1$ $85.9\%$
BHRF higuchi2024balanced $1.8$k $5$ $87.0\pm 0.4\%$
EF-adLIF (1 layer) $1.8$k $20$ $87.10\pm 0.46\%$
SE-adLIF (1 layer) $1.8$k $20$ $86.88\pm 0.40\%$
EF-adLIF (2 layers) $4.6$k $20$ $87.83\pm 0.51\%$
SE-adLIF (2 layers) $4.6$k $20$ $88.18\pm 0.36\%$
Table 1: Comparison of accuracies of different models on the test sets of SHD, SSC and ECG. The asterisk (*) of SHD* denotes that a fraction of the training data was used as a validation set. BHRF used $5\%$, whereas we used $20\%$ of the training data for validation. We report the test accuracy for the training epoch with the highest validation accuracy. For our model, SE-adLIF, we report the mean and standard deviation on $20$ different random initializations. Rec.: model contains recurrent synaptic connections; #Runs: number of runs with different random seeds to obtain the results. Test accuracy is provided as mean $\pm$ SD.
Our results show that recurrent adLIF networks clearly outperformed LIF network baselines on all tasks. For all considered datasets, recurrent SE-discretized adLIF networks performed better than previously considered recurrent SNNs. For SSC, their performance was slightly below that of the DCLS-model hammouamri2023learning, a feed-forward network using extensive delays trained via dilated convolutions. Networks composed of SE-discretized adLIF neurons (SE-adLIF networks) performed significantly better than those based on Euler-forward discretization (EF-adLIF networks) on SHD and SSC (significance values for a two-tailed t-test were $p<0.000001$ for SHD and $p<0.005$ for SSC). Small networks with a single recurrent layer on ECG performed on-par ($p=0.115$), while SE-adLIF networks significantly improved over EF-adLIF networks when larger networks with two recurrent layers were used ($p<0.02$). We found that EF-adLIF networks suffered from severe instabilities if neuron parameters were not constrained to values in which the decay rate exceeds the critical boundary of $r=1$, resulting in instabilities for example in the ECG task, see Section A in Appendix. The SE method is hence the preferred choice when adLIF neurons are used in a discretized form.
AdLIF neurons could, depending on their parameters, exhibit many different experimentally observed neuronal dynamics gerstner2014neuronal. We wondered whether networks trained on spatio-temporal classification tasks utilized the diverse dynamical behaviors of adLIF neurons. To that end, we investigated the resulting parameterizations of adLIF neurons in networks trained on SHD, and indeed found a heterogeneous landscape of neuron parameterizations, see Section B in the Appendix.
2.5 AdLIF networks accurately predict trajectories of dynamical systems
[…]
Note that in the spring-mass system the states are described by the displacement and velocity of the masses but only displacement information $\bm{x}[k]$ was available to the network. Hence, it is impossible to accurately predict the displacements of the masses at time $k+1$ from the displacements at time $k$ alone. The network must therefore learn to keep track of the longer-time dynamics of the system.
[…]
Additionally, we used this setup to compare the Symplectic-Euler discretization (SE-adLIF networks) with the Euler-Forward discretization (EF-adLIF networks). Since in this task the frequency bandwidth can be controlled directly via the spring coefficient, we generated spring-mass systems of increasing maximal frequency. We trained EF-adLIF networks and SE-adLIF networks under the same range of time-constants ($\tau_{u}$ and $\tau_{w}$) and a restricted range for the adaption parameter $a$. For Euler-Forward, $a$ was restricted between 00 and $a_{\mathrm{max}}$, where $a_{\mathrm{max}}$ is the maximal parameter value for $a$ that is stable under this discretization, resulting in a $[0,30]$ Hz range of frequencies that can be represented by the neurons for the chosen range of time-constants. For Symplectic-Euler, all frequencies below the Nyquist frequency are stable, so we simply choose $a_{\mathrm{max}}$ to achieve a frequency range of $[0,60]$ Hz. The results are shown in Fig. 4g. As expected, the two methods have similar performance at low frequencies. For dynamics with a larger frequency bandwidth however, EF-adLIF networks performed significantly worse. Additionally, the increased variance of the error indicates stability problems. These experimental results support our claim that the wider stability region of the SE-adLIF network allows the model to converge over a wide range of data frequencies.
2.6 Computational properties of adLIF networks
Our empirical results above demonstrate the superiority of oscillatory neuron dynamics over pure leaky integration in spiking neural networks, which is in line with prior studies bellec2018long; salaj2021spike; bittar2022surrogate; gangulySpikeFrequencyAdaptation2024a; higuchi2024balanced. In the following, we analyze the reasons behind this superiority.
Adaptation provides an inductive bias for temporal feature detection.
In gradient-based training of neural networks, the gradient determines to which features of the input a network ’tunes’ to. Hence, understanding how gradients depend on certain input features contributes to the understanding of network learning dynamics. When a recurrent SNN is trained with BPTT, the gradient propagates through the network via two different pathways: the recurrent synaptic connections and the implicit neuron-internal recurrence of the neuron state $s[k]$.
Figure 5: Inductive bias of adLIF gradients. a) Computational graph showing the state-to-state derivative $\frac{\partial s[t]}{\partial s[t-1]}$ back-propagating through time. $\mathbf{s}[k]$ denotes the state vector Eq. (5) b) Response of the membrane potential of a LIF neuron (left) and an adLIF neuron (right) to a single input spike. The shape of the derivative $\frac{\partial u[T]}{\partial u[t]}$ (bottom) matches the reversed impulse response function. c) Comparison of the derivative $\frac{\partial\mathcal{L}}{\partial\theta}$ for a wavelet input current. The multiplication from Eq. (12) of the input current with the state derivative is schematically illustrated for both, the adLIF and the LIF case. The frequency of the wavelet approximately matches the intrinsic frequency of the oscillation of the membrane potential oscillation of the adLIF neuron. The bar plot on the bottom shows the derivative $\frac{\partial\mathcal{L}}{\partial\theta}$ for both neurons, where color indicates input amplitude. d) Same as panel c but for a constant input current. e) Same as panel c but for different positions of the wavelet current. Middle plot shows the alignment between input and back-propagating derivative $\frac{\partial u[T]}{\partial u[t]}$ for the adLIF neuron. Input wavelet is given with a phase shift of 00,$-\frac{1}{2}P$ and $-\frac{3}{4}P$ with respect to the period $P$ of the adLIF neuron oscillation.
For the following analysis, we ignored the explicit recurrent synaptic connections and focused on how the backward gradient of the neuron state determines the magnitude of weight updates in different input scenarios, see Fig. 5a.
Consider the derivative $\frac{\partial u[T]}{\partial u[k]}$ of the membrane potential at a time step $T$ with respect to the membrane potential at some prior time step $k$. Intuitively, this derivative indicates how small perturbations of $u[k]$ influence $u[T]$. Fig. 5b shows this derivative for a LIF neuron (left) and an adLIF neuron (right). Because this derivative is the reverse of the model’s forward impulse response, it exhibits oscillations in the case of the adLIF neuron and reversed leaky integration for a LIF neuron. Consider a LIF or an adLIF neuron with a single synapse with weight $\theta$ and input $I[k]$. A loss signal $\frac{\partial L}{\partial u[T]}$ (set to $1$ in our illustrative example) is provided at time-step $T$. The resulting gradient $\frac{\partial L}{\partial\theta}$, used to compute the update of synaptic weight $\theta$, is given by
$\frac{\partial L}{\partial\theta}\propto\frac{\partial L}{\partial u[T]}\sum_{k=1}^{T}\frac{\partial u[T]}{\partial u[k]}I[k].$ (12)
This equation makes explicit that the weight change is proportional to the correlation between the input currents $I(T),I(T-1),I(T-2),\dots$ and the internal derivatives $\frac{\partial u[T]}{\partial u[T-1]},\frac{\partial u[T]}{\partial u[T-2]},\frac{\partial u[T]}{\partial u[T-3]},\dots$. This is illustrated in Fig. 5c for temporal input currents — realized as wavelets — at different amplitudes. The magnitudes of the resulting gradients for the LIF and adLIF model are quite complementary. The correlation between the wavelet and the oscillations of the adLIF neuron’s state-derivative results in a strongly amplitude-dependent gradient. In contrast, the leaky integration of the LIF neuron averages the positive and negative region in the input wave. The situation changes drastically for a constant input current, see Fig. 5d. For LIF neurons, the gradient $\frac{\partial L}{\partial\theta}$ strongly increases with increasing input current magnitude in this scenario. In contrast, the gradient of the adLIF neuron only weakly depends on the magnitude of the constant input current, in fact the gradient is nearly non-existent. This can be explained by the balance between positive and negative regions of the adLIF gradient (compare with Fig. 5b), resulting in almost zero if multiplied with a constant and summed over time.
The temporal sensitivity of the adLIF gradient is even more evident when we consider the gradients for different positions of a wavelet current (Fig. 5e). The sign and magnitude of the gradient strongly depend on the position of the wavelet for the adLIF neuron, but not for the LIF neuron. If the wavelet input is aligned with the oscillation of the back-propagating derivative, the resulting gradient is strongly positive. For a half-period ($-\frac{1}{2}P$) phase shift, the gradient is negative and for a $-\frac{3}{4}P$ phase shift, the resulting gradient is low in magnitude due to misalignment of oscillation and input. This gradient encourages the neuron to detect temporal features in the input, that is, temporally local changes in the input, either as changes in the spike rate (e.g. Fig. 2d) or in the input current (e.g. Fig. 5c,e) with specific timing. This sensitivity hence provides an inductive bias for spatio-temporal sequence processing tasks.
Networks of adLIF neurons tune to high-fidelity temporal features.
Through the rich dynamics and the consequential inductive bias towards learning temporal structure in the input, adLIF neurons should be well-suited for tasks in which spatio-temporal feature extraction is necessary. In order to investigate how well temporal input structure can be exploited by networks of adLIF neurons as compared to networks of LIF neurons, we considered a conceptual task that can be viewed as prototypical temporal pattern detection. We refer to this task as the burst sequence detection (BSD) task.
In the BSD task, a network has to classify temporal patterns of bursts from a population of $n$ input neurons, see Fig. 6a.
Figure 6: Temporal feature detection in adLIF networks. a) Two samples of classes $2$ and $17$ of the burst sequence detection (BSD) task (see main text). b) Classification error of adLIF and LIF networks with equal parameter count for different numbers of classes in the BSD task. c) Schematic illustration of network feature visualization. An initial noise sample $\bm{x_{0}}$ is passed through a trained network with frozen network parameters. The classification loss of the network output with respect to some predefined target class $c$ is computed and back-propagated through the network to obtain the gradient $\nabla_{X}L(X,c)|_{X^{0}}$ of the loss with respect to input $X^{0}$. This gradient is applied to the sample and the procedure is repeated to obtain a final sample $X^{K}$ after $K=400$ iterations. d) Samples generated by the feature visualization procedure from panel c from networks trained on the $20$-class BSD task. White dots denote the locations of the class-descriptive bursts. We generated samples for classes $2$ and $17$ which were the most misclassified classes of LIF and adLIF networks respectively. e) Samples generated from an adLIF network and a LIF network trained on SHD for different target classes $c$ (top) and the corresponding network output over time (bottom). The grey shaded area (at $t>100$) denotes the time span relevant for the loss, all outputs before this time span were ignored, see Methods for details.
This task is motivated from neuroscientific experiments which show the importance of sequences of transient increases of spike rates in cortex harvey2012choice; LISMAN199738. A class in this task is defined by a specific pre-defined temporal sequence of bursts across a fixed sub-population of three of these neurons. Other neurons emit a random burst at a random time each and additionally, neurons fire with a background rate of $50$ Hz. Bursts were implemented as smooth, transient increases of firing rate resulting in approximately $7$ spikes per burst, see Section 4.12 in Methods for details. We tested single-layer recurrent adLIF networks ($510$ neurons) and single-layer recurrent LIF networks with the same number of trainable parameters. AdLIF networks clearly outperformed LIF networks on this task, see Fig. 6b. For the case of $10$ classes, the adLIF network reached an average classification error of $2.31\%$ on the test set, whereas the LIF network only achieved a test error of $6.96\%$ despite a low training error ($<2\%$).
In order to evaluate to what extent the computations of the SNNs considered relied on temporal features of the input, we used a technique commonly applied to artificial neural networks that allows to visualize the input features that cause the network to predict a certain class erhan2009visualizing. The idea of this optimization-based feature visualization procedure is to generate an artificial data sample $X^{*}$ that maximally drives the network output towards a pre-defined target class $c$:
$X^{*}=\arg\min_{X}L(X,c),$ (13)
where $L(X,c)$ denotes the loss of the network output for input $X$ and target class $c$. In order to estimate $X^{*}$, one starts with a uniform noise input $X^{0}$ and updates the input using gradient descent to minimize the loss, i.e., the input $X^{k+1}$ after update ${k+1}$ is given by
$X^{k+1}=X^{k}-\frac{\eta}{\zeta^{k}}\nabla_{X}L(X,c)|_{X^{k}},$ (14)
where $\eta$ is the update step size, $\nabla_{X}f(X)|_{X^{k}}$ denotes the gradient of $f$ with respect to $X$ evaluated at $X^{k}$, and $\zeta_{k}$ is a normalization factor, see Fig. 6c and Section 4.13 in Methods for details. After each update, we applied additional regularization to the data sample $X^{k+1}$ (see Methods for details). We repeated this procedure for $K=400$ iterations, such that the final $X^{K}$ yielded a very strong prediction for class $c$.
We performed this feature visualization for the BSD task and for the SHD task. Fig. 6d shows the resulting artificial samples $X^{K}$ for an adLIF network and a LIF network trained on the BSD task. The class-defining burst timings are indicated as black circles with white filling.
The adLIF-generated samples exhibited a strong temporal structure that captures the relevant temporal structure of the class. This can be observed visually by comparing the position of strong activations with the class-defining burst timings. In contrast, the samples generated from the trained LIF network displayed less precise resemblance of class-descriptive features, and showed less temporal variation. This gap of specificity of the features in the generated samples of adLIF versus LIF networks might explain the performance gap between these: While the less precise temporal tuning of LIF networks suffices to achieve a high accuracy on the training data, it falls behind in terms of generalization on the test set, due to confusion of temporal features from different classes. Interestingly, the same analysis for an ALIF network yinAccurateEfficientTimedomain2021 with threshold adaptation revealed that the temporal tuning of ALIF networks is comparable to that of LIF networks, indicating the importance of oscillatory dynamics for temporal feature detection, see Section D in the Appendix.
Similar results were obtained for SHD, Fig. 6e, where we applied $K=200$ iterations. The underlying class-descriptive features in the SHD task are less clearly visible, since it is instantiated from natural speech recordings. Nevertheless, one can clearly observe richer temporal structure of the adLIF-generated samples. These samples indicate, that the two network models tune to very different features of the input, which is again in alignment with the above reported inductive bias of the gradient. While LIF networks tend to tune to certain spike rates of different input neurons over prolonged durations, adLIF neurons rather tune to local variations of the spike rates. In summary, our analysis supports the hypothesis that the superior performance of SNNs based on adLIF neurons on datasets like SHD stems from the fact that temporal features can effectively be learned and detected.
Inherent normalization properties of adLIF neurons.
Artificial neural networks as well as spiking neural networks are often trained using normalization techniques that normalize the input to the network layers batchnorm. For SNNs, several normalization techniques have been developed kim2021revisitingbatchnormalizationtraining; vicente-solaSpikingNeuralNetworks2023. These techniques are however problematic from an implementation perspective, in particular when such networks are implemented in neuromorphic hardware. In batch normalization through time kim2021revisitingbatchnormalizationtraining for example, a different normalization is applied to the membrane potential at every time step. In contrast, all results reported in this article have been achieved without an explicit normalization technique, indicating that such normalization is not necessary for networks of adLIF neurons.
We argue that good performance without normalization is possible due to the negative feedback loop through the adaption current in adLIF neurons (Eqs. (2), (3)) which inherently stabilizes neuron responses as long as neurons are in the stable regime. In addition, the oscillatory sub-threshold response (Fig. 1b) tends to filter out constant offsets in the forward pass. Similarly, the oscillating gradient (Fig. 5b) tends to filter out constant activation offsets during training, thus stabilizing training.
To test the stabilizing effect of adLIF neurons, we investigated how a constant offset in the input during test time effects the accuracy of RSNNs. To this end, tested LIF and adLIF networks trained on the SHD dataset (Table 1) on biased SHD test examples which we obtained by adding a constant offset to each input dimension. More precisely, the biased sample $\hat{\bm{x}}[k]$ at time step $k$ was given by $\hat{\bm{x}}[k]=\bm{x}[k]+\kappa\bar{x}$, where $\bar{x}$ is the mean over all input dimensions and time steps and $\kappa\geq 0$ scales the bias strength. The classification accuracy of the network on these biased test samples as a function of the bias coefficient $\kappa$ is shown in Fig. 7a.
[…]
Surprisingly, even for large bias values the adLIF network maintained a high classification accuracy. In comparison, the accuracy of a LIF network dropped rapidly with increased $\kappa$.
In a second experiment, instead of a constant bias, we added background noise via random spikes to the raw input data. Again, the adLIF network was surprisingly robust to this noise even for high background noise rates robustness between adLIF and LIF networks, see Fig. 7b,c.
These results support the claim that normalization methods are not necessary to train noise-robust high-performance adLIF networks, which represents a substantial advantage of this model over LIF-based SNNs in neuromorphic applications.
3 Discussion
Spiking neural network models are the basis of many neuromorphic systems. For a long time, these networks used leaky integrate-and-fire neurons as their fundamental computational units. More recent work has shown that networks of adaptive spiking neurons outperform LIF networks in spatio-temporal processing tasks. However, a deep understanding of the mechanisms that underlie their superiority was lacking. In this article, we investigated the underpinnings of the computational capabilities of networks of adaptive spiking neurons.
We first showed how stability issues of adLIF neurons discretized with the Euler-Forward method can be avoided with the Symplectic-Euler method geometric_numerical_integration, resulting in the SE-adLIF model. We analytically derived its favorable stability properties and demonstrated its practical superiority on several benchmark datasets. Many digital neuromorphic chips implement time-discretized SNNs merolla2014million; furber2014spinnaker; orchardEfficientNeuromorphicSignal2021; frenkel2022reckon. Adaptive neurons are an attractive model for such systems as they only add a single additional state variable per neuron. As synaptic connections are typically dominating implementation costs, this approximate doubling of resources needed for neuron dynamics is well justified by the improved performance. For example, Fig. 4f shows that there exist tasks where adLIF networks can achieve superior performance to LIF networks with orders of magnitudes less parameters. Compared to the standard Euler-discretized adLIF model, the computational operations needed to implement the SE-adLIF model are identical. Hence, the improved stability properties of this model are practically for free.
Our analysis in Section Computational properties of adLIF networks indicates that adLIF should be well-suited to learn relevant temporal features from input sequences. Interestingly, our investigations revealed that LIF networks are surprisingly weak in that respect. This is witnessed by its low performance at the burst sequence detection task (Fig. 6) as well as by our input-feature analysis of trained LIF networks (Fig. 6 panels d and f). This is surprising, as theoretical arguments suggest that SNNs in general should be efficient in temporal computing tasks maass1997networks. Our analysis suggests that the gradients in LIF networks fail to detect such temporal features, while for adLIF networks, these gradients are actually biased towards those. Our empirical evaluation supports this claim as adLIF networks excel at the burst sequence detection task (Fig. 6), a task which we designed specifically to test temporal feature detection capabilities of SNNs (Fig. 6). The input-feature analysis for trained adLIF networks further supports this view (Fig. 6 panels b and d).
The auto-regressive task on a complex spring-mass system (Fig. 4) can also be seen as a conceptual task designed to investigate the capabilities of SNNs to predict the behavior of complex oscillatory systems. Note that, although the trajectories of masses that have to be predicted are periodic, the period is very long due to the complex interactions of the four masses. This conceptual task is of high relevance as oscillations are ubiquitous in physical systems and biology, for example in limb movement patterns kilnerHumanCorticalMuscle2000; sanesOscillationsLocalField1993. The superiority of adLIF networks in this task can be explained by the principles of physics-informed neural networks raissiPhysicsinformedNeuralNetworks2019. In this framework, underlying physical laws of some training data are molded into the architecture of neural networks, such that the functions learned by these networks naturally follow these laws. Obviously, the oscillatory sub-threshold behavior of adLIF neurons fits well to the oscillatory dynamics of the spring-mass system.
Finally, we empirically demonstrated the robustness of adLIF networks towards perturbations in the input, showcasing their invariance to shifts in the mean input strength. Surprisingly, the accuracy of adLIF networks on the SHD dataset maintained a high level ($>80\%$) even after doubling the mean input strength, whereas LIF networks drop to an accuracy of $10\%$ already for a much lower increase. We argue that this inheritance alleviates the need for layer normalization techniques in other types of artificial and spiking neural networks. Indeed, all our results were achieved without explicit normalization techniques. This finding is in particular relevant for neuromorphic implementations of SNNs, as explicit normalization is hard to implement in neuromorphic hardware, in particular for recurrent SNNs.
Oscillatory neural network dynamics have been studied not only in the context of SNNs. Several works rusch2021coupled; dubininFadingMemoryInductive2024b; effenbergerFunctionalRoleOscillatory2023b identified favorable properties adopted by models utilizing some form of oscillations. Rusch et al. rusch2021coupled for example studied an RNN architecture in which recurrent dynamics were given by the equation of motion of a damped harmonic oscillator. The authors found, that these oscillations not only alleviate the vanishing and exploding gradient problem (glorot), but also perform well on a large variety of benchmarks. Effenberger et al. effenbergerFunctionalRoleOscillatory2023b proposed oscillating networks as a model for cortical columns. In the field of artificial neural networks, the recent advent of state space models voelkerLegendreMemoryUnits2019b; guMambaLinearTimeSequence2023a; guptaDiagonalStateSpaces2022; smithSimplifiedStateSpace2023a and linear recurrent networks orvietoResurrectingRecurrentNeural2023a introduces a paradigm shift in sequence processing, where information is transported through constrained linear state transitions instead of being recurrently propagated between nonlinear neurons, as was previously the case in traditional recurrent neural networks elmanFindingStructureTime1990. Similar to SNNs, state space models are obtained by discretizing continuous ordinary differential equations (ODEs) to form recurrent neural networks. Although spiking neuron models have always been derived from discretizing differential equations to obtain recurrent linear state transitions maass2001pulsed, earlier neuron models, such as the LIF neuron, lack the temporal dynamics necessary to effectively propagate time-sensitive information. The relation between SNNs and state space models has recently been discussed balRethinkingSpikingNeural2024; schoneScalableEventbyeventProcessing2024a.
In summary, we have shown that networks of adaptive LIF neurons provide a powerful model of computation for neuromorphic systems. Stability issues during training can provably be avoided by the use of a suitable discretization method. Our results indicate that the properties of these neurons, in particular their sub-threshold oscillatory response, provide the basis for their spatio-temporal processing capabilities.
4 Methods
4.1 Details to simulations in Figure 2
For plots in Fig. 2b-c, we used the adLIF neuron parameters $\tau_{w}=60$ ms, $\tau_{u}=15$ ms, $a=120$, $\vartheta=\infty$. For Fig. 2d-f the adLIF neuron parameters were $\tau_{w}=200$ ms, $\tau_{u}=125$ ms, $a=100$, $\vartheta=\infty$. For Fig. 2a and g, parameters for the LIF neuron were $\tau_{u}=125$ ms and $\vartheta=\infty$.
The spike trains for Fig. 2d and e were generated deterministically in the following way. We first computed a spike rate $s[t]\in[0,0.2]$ for each time step $t$ according to $s[t]=0.2(0.5+0.5\sin\left(2\pi tF\Delta t\right))$, where $F$ is the oscillation frequency of the sinus signal in Hz ($F=10$ Hz for Fig. 2d, $F=7$ Hz for Fig. 2e), $\Delta t=1$ ms the sampling time step, and $t\in[0,...,T]$. Then, the spike train $S[t]$ was computed by cumulatively summing over the spike rates in an integrate-and-fire manner with $v[t]=v[t-1]+s[t]-S[t]$, where $S[t]=\Theta(v[t]-1)$ with Heaviside step function $\Theta$ and $v[0]=0$.
4.2 Derivation of matrices $\bar{A}_{\mathrm{SE}}$ and $\bar{B}_{\mathrm{SE}}$ for the SE-discretized adLIF neuron
Here, we provide the derivation of the matrices $\bar{A}_{\mathrm{SE}}$ and $\bar{B}_{\mathrm{SE}}$ for the SE-discretized adLIF neuron given in Eq. (11). To rewrite the state update equations of the SE-adLIF neuron model, given by
$\hat{u}[k]$ $=\alpha u[k-1]+(1-\alpha)\left(-w[k-1]+I[k]\right)$ (15a)
$w[k]$ $=\beta w[k-1]+(1-\beta)(au[k]+bS[k]),$ (15b)
into the canonical state-space representation
$\bm{s}[k]=\begin{pmatrix}u[k]\\ w[k]\end{pmatrix}=\bar{A}_{\mathrm{SE}}\bm{s}[k-1]+\bar{B}_{\mathrm{SE}}\bm{x}[k],$ (16)
we substitute $u[k]$ in Eq. (15b) by $\hat{u}[k]$ from Eq. (15a). Since we study the sub-threshold dynamics of the neuron (assuming $S[k]=0$), we can substitute $\hat{u}[k]$ by $u[k]$, since $u[k]=\hat{u}[k]\cdot(1-S[k])$. This yields the update equation for $w[k]$ given by
$w[k]=\beta w[k-1]+(1-\beta)\left(a\underbrace{\left(\alpha u[k-1]+(1-\alpha)\left(-w[k-1]+I[k]\right)\right)}_{\hat{u}[k]}+bS[k]\right).$ (17)
Since this formulation gives the new state of adaptation variable $w[k]$ as function of the previous states $u[k-1]$ and $w[k-1]$, it can be transformed into a matrix formulation
$\begin{pmatrix}u[k]\\ w[k]\end{pmatrix}=\underbrace{\begin{pmatrix}\alpha&-(1-\alpha)\\ a(1-\beta)\alpha&\quad\beta-a(1-\beta)(1-\alpha)\\ \end{pmatrix}}_{\bar{A}_{\mathrm{SE}}}\begin{pmatrix}u[k-1]\\ w[k-1]\end{pmatrix}+\underbrace{\begin{pmatrix}(1-\alpha)&0\\ 0&b(1-\beta)\\ \end{pmatrix}}_{\bar{B}_{\mathrm{SE}}}\begin{pmatrix}I[k]\\ S[k]\end{pmatrix}.$ (18)
4.3 Proof of stability bounds for the continuous adLIF model
For all following analyses in Sections 4.3 to 4.7, we consider the subthreshold regime, i.e., we assume a spike threshold $\vartheta=\infty$ such that $S[k]=0$ for all $k$, and we assume no external inputs $I$.
In this section, we prove that the continuous-time adLIF neuron exhibits stable sub-threshold dynamics for all $a>-1$.
Lemma 4.1.
The continuous adLIF neuron model from Eqs. (2) and (3) is stable in the sub-threshold regime for all $\tau_{u}>0,\tau_{w}>0,a>-1$.
Proof.
In general, a continuous-time linear dynamical system $\dot{x}=Ax$ is Lyapunov-stable, if the real parts $\Re(\lambda_{1,2})$ of both eigenvalues $\lambda_{1,2}$ of matrix $A$ satisfy $\Re(\lambda_{1,2})\leq 0$ linear-systems-theory. For the adLIF model, the matrix $A$ is given by
$A=\begin{pmatrix}-\frac{1}{\tau_{u}}&-\frac{1}{\tau_{u}}\\ \frac{a}{\tau_{w}}&-\frac{1}{\tau_{w}}\end{pmatrix},$ (19)
with eigenvalues
$\lambda_{1}$ $=\frac{-\tau_{u}-\tau_{w}+\sqrt{-4a\tau_{u}\tau_{w}+(\tau_{u}-\tau_{w})^{2}}}{2\tau_{u}\tau_{w}}$ (20)
$\lambda_{2}$ $=\frac{-\tau_{u}-\tau_{w}-\sqrt{-4a\tau_{u}\tau_{w}+(\tau_{u}-\tau_{w})^{2}}}{2\tau_{u}\tau_{w}}.$ (21)
In the complex-valued case (where the discriminant $-4a\tau_{u}\tau_{w}+(\tau_{u}-\tau_{w})^{2}<0$), the real part $\Re(\lambda_{1,2})$ is given by
$\Re(\lambda_{1,2})=\frac{-\tau_{u}-\tau_{w}}{2\tau_{u}\tau_{w}},$ (22)
where $\Re(\lambda_{1,2})<0$ since $\tau_{u},\tau_{w}>0$. In the case of real eigenvalues, $\lambda_{2}<0$ is always true (since Eq. (21) only contains negative terms), whereas $\lambda_{1}\leq 0$ is only true if
$\sqrt{-4a\tau_{u}\tau_{w}+(\tau_{u}-\tau_{w})^{2}}\leq\tau_{u}+\tau_{w}$ (23)
$a\geq-\frac{(\tau_{u}+\tau_{w})^{2}-(\tau_{u}-\tau_{w})^{2}}{4\tau_{u}\tau_{w}}$ (24)
Hence, the continuous-time adLIF neuron is stable for all
$a\geq-1.$ (25)
This proves, that for $a\geq-1$, the continuous-time adLIF neuron is Lyapunov stable. ∎
4.4 Derivation of stability bounds for EF-adLIF
In this section we derive the stability bounds of the EF-adLIF neuron with respect to parameter $a$. These bounds provide the basis for the subsequent proofs. The state update equations of the Euler-Forward discretized adLIF neuron are given by
$\dot{\bm{s}}(t)=\begin{pmatrix}\dot{u}(t)\\ \dot{w}(t)\end{pmatrix}=\bar{A}_{\mathrm{EF}}\bm{s}(t)+\bar{B}_{\mathrm{EF}}\bm{x}(t)$
with
$\bar{A}_{\mathrm{EF}}=\begin{pmatrix}\alpha&-(1-\alpha)\\ a(1-\beta)&\beta\end{pmatrix}\hskip 30.0pt\bar{B}_{\mathrm{EF}}=\begin{pmatrix}(1-\alpha)&0\\ 0&b(1-\beta)\\ \end{pmatrix}$ (26)
with $\alpha=e^{-\frac{\Delta t}{\tau_{u}}}$ and $\beta=e^{-\frac{\Delta t}{\tau_{w}}}$. This system is asymptotically stable, if the spectral radius $\rho$, given by $\rho=\max\left(|\lambda_{1}|,|\lambda_{2}|\right)$ is less than $1$. We differentiate between two cases: complex-valued and real-valued eigenvalues. We assume given time constants $\tau_{u},\tau_{w}>0$ and calculate the bound as function of parameter $a$. To simplify the notation, we introduce $\bar{\alpha}=(1-\alpha)$ and $\bar{\beta}=(1-\beta)$.
Lemma 4.2.
The EF-adLIF model is asymptotically stable and oscillating in the sub-threshold regime, for $a\in~{}(a_{0}^{\text{EF}},a_{\mathrm{max}}^{\text{EF}})$, with $a_{0}^{\text{EF}}=\frac{(\alpha-\beta)^{2}}{4(1-\alpha)(1-\beta)}$ and $a_{\mathrm{max}}^{\text{EF}}=\frac{1-\alpha\beta}{(1-\alpha)(1-\beta)}$. The real part of the eigenvalues is strictly positive and given by $\Re(\lambda_{1,2})=\frac{\alpha+\beta}{2}>0$.
Proof.
The characteristic polynomial $\chi(\bar{A}_{\mathrm{EF}})$ is given by
$\chi(\bar{A}_{\mathrm{EF}})=\lambda^{2}+\lambda\left(-\alpha-\beta\right)+\beta\alpha+a\bar{\alpha}\bar{\beta},$ (27)
with discriminant
$\Delta_{\lambda}$ $=(\alpha-\beta)^{2}-4a\bar{\alpha}\bar{\beta}.$ (28)
$\chi(\bar{A}_{\mathrm{EF}})$ admits complex solutions for $\Delta_{\lambda}<0$ depending on $a$, with
$a$ $>\frac{(\alpha-\beta)^{2}}{4\bar{\alpha}\bar{\beta}}=a_{0}^{\text{EF}},$ (29)
Complex roots of $\chi(\bar{A}_{\mathrm{EF}})$ yield the complex-conjugate eigenvalues
$\lambda_{1,2}=\frac{\alpha+\beta}{2}\pm i\frac{\sqrt{-\Delta_{\lambda}}}{2}.$ (30)
This proves that the complex eigenvalues have a strictly positive real part given by $\Re(\lambda_{1,2})=\frac{\alpha+\beta}{2}>0$. In that case the spectral radius $\rho$ is defined as $\rho(\bar{A}_{\mathrm{EF}})=|\lambda_{1}|=|\lambda_{2}|$, where $|z|=\sqrt{\mathrm{Re}(z)^{2}+\mathrm{Im}(z)^{2}}$ is the modulus of the complex number with
$\rho(\bar{A}_{\mathrm{EF}})$ $=\frac{1}{2}\sqrt{\left(\beta+\alpha\right)^{2}-\Delta_{\lambda}}$ (31)
$=\frac{1}{2}\sqrt{\left(\beta+\alpha\right)^{2}-\left(\beta-\alpha\right)^{2}+4\bar{\alpha}}\bar{\beta}$ (32)
$=\sqrt{\alpha\beta+a\bar{\alpha}\bar{\beta}}.$ (33)
In the complex regime, the system is stable when $\rho(\bar{A}_{\mathrm{EF}})<1$, and thus,
$a<\frac{1-\alpha\beta}{\bar{\alpha}\bar{\beta}}=a_{\mathrm{max}}^{\text{EF}},$ (34)
where $a_{\mathrm{max}}^{\text{EF}}$ gives the upper stability bound.
The system is asymptotically stable in the sub-threshold regime for $a\in~{}(\frac{(\alpha-\beta)^{2}}{4\bar{\alpha}\bar{\beta}},\frac{1-\alpha\beta}{\bar{\alpha}\bar{\beta}})$. In this bound, the system has complex eigenvalues and thus admits oscillations. ∎
Lemma 4.3.
EF-adLIF is asymptotically stable and not oscillating in the sub-threshold regime for $a\in(-1,a_{0}^{\text{EF}}]$ with $a_{0}^{\text{EF}}=\frac{(\alpha-\beta)^{2}}{4\bar{\alpha}\bar{\beta}}$.
Proof.
The characteristic polynomial $\chi(\bar{A}_{\mathrm{EF}})$ is given by
$\chi(\bar{A}_{\mathrm{EF}})=\lambda^{2}+\lambda\left(-\alpha-\beta\right)+\beta\alpha+a\bar{\alpha}\bar{\beta},$ (35)
with discriminant
$\Delta_{\lambda}$ $=(\alpha-\beta)^{2}-4a\bar{\alpha}\bar{\beta}$ (36)
$\chi(\bar{A}_{\mathrm{EF}})$ admits real solutions for $\Delta_{\lambda}\geq 0$ depending on $a$, with
$\Delta_{\lambda}$ $\geq 0$ (37)
$(\alpha-\beta)^{2}$ $\geq 4a\bar{\alpha}\bar{\beta}$ (38)
$a$ $\leq\frac{(\alpha-\beta)^{2}}{4\bar{\alpha}\bar{\beta}}=a_{0}^{\text{EF}},$ (39)
$\chi(\bar{A}_{\mathrm{EF}})$ thus admits real solutions in the interval $\left(\infty,a_{0}^{\text{EF}}\right]$ with real eigenvalues
$\lambda_{1,2}=\frac{\alpha+\beta\pm\sqrt{\Delta_{\lambda}}}{2}.$ (40)
For $a=a_{0}^{\text{EF}}$, the singular real root of $\chi(\bar{A}_{\mathrm{EF}})$ is
$\lambda_{0}=\frac{\alpha+\beta}{2}<1.$ (41)
For lower values of $a$ we have the following asymptotic behavior,
$\lim_{a\to-\infty}\lambda_{1}$ $=\frac{\alpha+\beta+\sqrt{\Delta_{\lambda}}}{2}=\infty$ (42)
$\lim_{a\to-\infty}\lambda_{2}$ $=\frac{\alpha+\beta-\sqrt{\Delta_{\lambda}}}{2}=-\infty,$ (43)
where $\lambda_{1}$ is the term with the higher absolute value and hence determines the spectral radius $\rho(\bar{A}_{\mathrm{EF}})$. It exists a value $a_{\mathrm{min}}^{\text{EF}}$ such that $\lambda_{1}(a_{\mathrm{min}}^{\text{EF}})=1$,
$\lambda_{1}$ $=\frac{\alpha+\beta+\sqrt{(\alpha-\beta)^{2}-4a\bar{\alpha}\bar{\beta}}}{2}=1,$ (44)
$a_{\mathrm{min}}^{\text{EF}}$ $=\frac{1-\alpha-\beta+\alpha\beta}{-\bar{\alpha}\bar{\beta}}=\frac{\bar{\alpha}\bar{\beta}}{-\bar{\alpha}\bar{\beta}}=-1,$ (45)
and the system is unstable for $a<a_{\mathrm{min}}^{\text{EF}}=-1$. The system is thus asymptotically stable in the sub-threshold regime for $a\in(-1,a_{0}^{\text{EF}}]$. In this bound, the system has real eigenvalues and thus doesn’t admit oscillations. ∎
Corollary 4.3.1.
In the sub-threshold regime (i.e. $\vartheta=\infty$), for $\tau_{u},\tau_{w},\Delta t>0$, $\alpha=e^{-\frac{\Delta t}{\tau_{u}}}$ and $\beta=e^{-\frac{\Delta t}{\tau_{w}}}$, EF-adLIF is asymptotically stable for $a\in(-1,a_{\mathrm{max}}^{\text{EF}})$ with $a_{\mathrm{max}}^{\text{EF}}=\frac{1-\alpha\beta}{(1-\beta)(1-\alpha)}$.
This is a consequence of Lemma 4.2 and Lemma 4.3.
4.5 Derivation of the stability bounds of SE-adLIF
Analogously to Section 4.4, we can compute the stability bounds for the SE-discretized adLIF neuron (SE-adLIF) with respect to parameters $\tau_{u}$, $\tau_{w}$, and $a$. Again, we assume time constants $\tau_{u},\tau_{w}>0$ as given and calculate the stability bounds with respect to parameter $a$. Recall state transition matrix $\bar{A}_{\mathrm{SE}}$ from Eq. (11):
$\bar{A}_{\mathrm{SE}}=\begin{pmatrix}\alpha&-(1-\alpha)\\ a(1-\beta)\alpha&\quad\beta-a(1-\beta)(1-\alpha)\\ \end{pmatrix}$ (46)
with $\alpha=e^{-\frac{\Delta t}{\tau_{u}}}$ and $\beta=e^{-\frac{\Delta t}{\tau_{w}}}$. To simplify notation, we again introduce $\bar{\alpha}=(1-\alpha)$ and $\bar{\beta}=(1-\beta)$.
Lemma 4.4.
SE-adLIF is stable and oscillating in the sub-threshold regime for $a\in(a_{1}^{\text{SE}},a_{2}^{\text{SE}})$, with $a_{1}^{\text{SE}}=\frac{\left(\sqrt{\beta}-\sqrt{\alpha}\right)^{2}}{\bar{\beta}\bar{\alpha}}$ and $a_{2}^{\text{SE}}=\frac{\left(\sqrt{\beta}+\sqrt{\alpha}\right)^{2}}{\bar{\beta}\bar{\alpha}}$. The spectral radius is independent of $a$ and given by $\rho=\sqrt{\beta\alpha}<1$.
Proof.
The characteristic polynomial of $\chi(\bar{A}_{\mathrm{SE}})$ is
$\chi(\bar{A}_{\mathrm{SE}})=\lambda^{2}+\lambda\left(\bar{\beta}\bar{\alpha}a-\beta-\alpha\right)+\beta\alpha,$ (47)
with discriminant $\Delta_{\lambda}$ given by
$\Delta_{\lambda}$ $=\left(\bar{\beta}\bar{\alpha}a-\beta-\alpha\right)^{2}-4\beta\alpha$
$=\bar{\beta}\bar{\alpha}\left[\bar{\beta}\bar{\alpha}a^{2}-2(\beta+\alpha)a+\frac{(\beta-\alpha)^{2}}{\bar{\beta}\bar{\alpha}}\right].$ (48)
$\chi(\bar{A}_{\mathrm{SE}})$ gives complex solutions for $\Delta_{\lambda}<0$, which depends on the range of negative values of the polynomial part $p(a)$ of $\Delta_{\lambda}$, given by
$p(a)=\bar{\beta}\bar{\alpha}a^{2}-2(\beta+\alpha)a+\frac{(\beta-\alpha)^{2}}{\bar{\beta}\bar{\alpha}}.$ (49)
From this polynomial, we can again compute a discriminant $\Delta_{a}$ as
$\Delta_{a}=16\beta\alpha.$ (50)
Since $\Delta_{a}>0$ the roots of this polynomial are given by two values $a_{1}^{\text{SE}}$ and $a_{2}^{\text{SE}}$ according to
$a_{1}^{\text{SE}}=\frac{\left(\sqrt{\beta}-\sqrt{\alpha}\right)^{2}}{\bar{\beta}\bar{\alpha}},\quad a_{2}^{\text{SE}}=\frac{\left(\sqrt{\beta}+\sqrt{\alpha}\right)^{2}}{\bar{\beta}\bar{\alpha}}.$ (51)
As the coefficient $\bar{\beta}\bar{\alpha}$ in Eq. (48) is always positive, $\Delta_{\lambda}$ is negative when $a_{1}^{\text{SE}}<a<a_{2}^{\text{SE}}$ resulting in complex-conjugate eigenvalues $\lambda_{1,2}$, given by the roots of $\chi(\bar{A}_{\mathrm{SE}})$, implying oscillatory behavior of the membrane potential. These eigenvalues are given by
$\lambda_{1,2}=-\frac{\left(\bar{\beta}\bar{\alpha}a-\beta-\alpha\right)}{2}\pm i\frac{\sqrt{-\Delta_{\lambda}}}{2}.$ (52)
In that case the spectral radius is defined as $\rho(A)=|\lambda_{1}|=|\lambda_{2}|$, where $|z|=\sqrt{\mathrm{Re}(z)^{2}+\mathrm{Im}(z)^{2}}$ is the modulus of the complex eigenvalues, such that
$r=\rho(A)$ $=\frac{1}{2}\sqrt{\left(\bar{\beta}\bar{\alpha}a-\beta-\alpha\right)^{2}-\Delta_{\lambda}}$ (53)
$=\frac{1}{2}\sqrt{\left(\bar{\beta}\bar{\alpha}a-\beta-\alpha\right)^{2}-\left(\bar{\beta}\bar{\alpha}a-\beta-\alpha\right)^{2}+4\beta\alpha}$ (54)
$=\sqrt{\beta\alpha}.$ (55)
Hence, the spectral radius, which we also refer to as the decay rate $r$ in the main text, is $\sqrt{\beta\alpha}$ which is always strictly less than $1$ due to $\beta,\alpha\in\left(0,1\right)$. Therefore, the SE-adLIF neuron is stable over the entire range of parameters (provided $\tau_{u},\tau_{w}>0$) where the matrix $\bar{A}_{\mathrm{SE}}$ exhibits complex eigenvalues. ∎
Lemma 4.5.
SE-adLIF is stable and not oscillating in the sub-threshold regime for $a\in(-1,a_{1}^{\text{SE}}]\cup[a_{2}^{\text{SE}},a_{\mathrm{max}}^{\text{SE}})$ with $a_{\mathrm{max}}^{\text{SE}}=\frac{(1+\beta)(1+\alpha)}{(1-\beta)(1-\alpha)}$ and $a_{1}^{\text{SE}}$, $a_{2}^{\text{SE}}$ as defined in Lemma 4.4.
Proof.
The characteristic polynomial of $\chi(\bar{A}_{\mathrm{SE}})$ is
$\chi(\bar{A}_{\mathrm{SE}})=\lambda^{2}+\lambda\left(\bar{\beta}\bar{\alpha}a-\beta-\alpha\right)+\beta\alpha,$ (56)
with discriminant $\Delta_{\lambda}$ given by
$\Delta_{\lambda}$ $=\left(\bar{\beta}\bar{\alpha}a-\beta-\alpha\right)^{2}-4\beta\alpha$
$=\bar{\beta}\bar{\alpha}\left[\bar{\beta}\bar{\alpha}a^{2}-2(\beta+\alpha)a+\frac{(\beta-\alpha)^{2}}{\bar{\beta}\bar{\alpha}}\right].$ (57)
The derivations from Lemma 4.4 imply that $\chi(\bar{A}_{\mathrm{SE}})$ admits real solutions in the intervals $\left(-\infty,a_{1}^{\text{SE}}\right]$ and $\left[a_{2}^{\text{SE}},\infty\right)$, yielding real-valued eigenvalues,
$\lambda_{1,2}=\frac{-\left(\bar{\beta}\bar{\alpha}a-\beta-\alpha\right)\pm\sqrt{\Delta_{\lambda}}}{2}.$ (58)
For $a=a_{1}^{\text{SE}}$ and $a=a_{2}^{\text{SE}}$, $\Delta_{\lambda}=0$ and we have two singular solutions,
$\lambda^{0}_{1}$ $=-\frac{\left(\bar{\beta}\bar{\alpha}a_{1}^{\text{SE}}-\beta-\alpha\right)}{2}=\sqrt{\beta\alpha}$ (59)
$\lambda^{0}_{2}$ $=-\frac{\left(\bar{\beta}\bar{\alpha}a_{2}^{\text{SE}}-\beta-\alpha\right)}{2}=-\sqrt{\beta\alpha}.$ (60)
For both singular solutions the spectral radius is $\rho(\bar{A}_{\mathrm{SE}})=|r^{0}_{1}|=|r^{0}_{2}|=\sqrt{\beta\alpha}$ which is less than one, resulting in asymptotic stability.
In order to study the stability for $a<a_{1}^{\text{SE}}$ and $a>a_{2}^{\text{SE}}$ we need to determine the asymptotic behavior of $\lambda_{1,2}$ for $a\to-\infty$ and $a\to\infty$, as it allows us to determine which eigenvalue will constrain the stability of the system. In the following propositions, we prove the stability of the system in the intervals $(-1,a_{1}^{\text{SE}}]$ and $[a_{2}^{\text{SE}},a_{\mathrm{max}}^{\text{SE}})$ independently.
Proposition 1.
SE-adLIF is asymptotically stable for $a\in(-1,a_{1}^{\text{SE}}]$.
For $a<a_{1}^{\text{SE}}$, $\rho(\bar{A}_{\mathrm{SE}})$ is determined by $|\lambda_{1}|$ as for $a\to-\infty$, we have
$\lim_{a\to-\infty}\lambda_{1}$ $=\lim_{a\to-\infty}\frac{-\left(\bar{\beta}\bar{\alpha}a-\beta-\alpha\right)+\sqrt{\left(\bar{\beta}\bar{\alpha}a-\beta-\alpha\right)^{2}-4\beta\alpha}}{2}=\infty$ (61)
$\lim_{a\to-\infty}\lambda_{2}$ $=\lim_{a\to-\infty}\frac{-\left(\bar{\beta}\bar{\alpha}a-\beta-\alpha\right)-\sqrt{\left(\bar{\beta}\bar{\alpha}a-\beta-\alpha\right)^{2}-4\beta\alpha}}{2}=0.$ (62)
Thus $\lambda^{0}_{1}<1<\lim\limits_{a\to-\infty}\lambda_{1}$. We can find a value $a_{\mathrm{min}}^{\text{SE}}$ such that $\lambda_{1}(a_{\mathrm{min}}^{\text{SE}})=1$, given by
$\lambda_{1}$ $=\frac{-\left(\bar{\beta}\bar{\alpha}a_{\mathrm{min}}^{\text{SE}}-\beta-\alpha\right)+\sqrt{\left(\bar{\beta}\bar{\alpha}a_{\mathrm{min}}^{\text{SE}}-\beta-\alpha\right)^{2}-4\beta\alpha}}{2}=1$ (63)
$a_{\mathrm{min}}^{\text{SE}}$ $=\frac{-\beta\alpha+\beta+\alpha-1}{\bar{\beta}\bar{\alpha}}=\frac{-\bar{\beta}\bar{\alpha}}{\bar{\beta}\bar{\alpha}}=-1.$ (64)
Proposition 2.
SE-adLIF is asymptotically stable for $a\in[a_{2}^{\text{SE}},a_{\mathrm{max}}^{\text{SE}})$ with $a_{\mathrm{max}}^{\text{SE}}=\frac{(1+\beta)(1+\alpha)}{(1-\beta)(1-\alpha)}$.
For $a>a_{2}^{\text{SE}}$, $\rho(\bar{A}_{\mathrm{SE}})$ is determined by $|\lambda_{2}|$ as for $a\to\infty$, we have
$\lim_{a\to\infty}\lambda_{1}$ $=\lim_{a\to\infty}\frac{-\left(\bar{\beta}\bar{\alpha}a-\beta-\alpha\right)+\sqrt{\left(\bar{\beta}\bar{\alpha}a-\beta-\alpha\right)^{2}-4\beta\alpha}}{2}=0$ (65)
$\lim_{a\to\infty}\lambda_{2}$ $=\lim_{a\to\infty}\frac{-\left(\bar{\beta}\bar{\alpha}a-\beta-\alpha\right)-\sqrt{\left(\bar{\beta}\bar{\alpha}a-\beta-\alpha\right)^{2}-4\beta\alpha}}{2}=-\infty.$ (66)
Thus $\lim\limits_{a\to\infty}\lambda_{2}<-1<\lambda^{0}_{2}$. We can find a value $a_{\mathrm{max}}^{\text{SE}}$ such that $\lambda_{2}(a_{\mathrm{max}}^{\text{SE}})=-1$, given by
$\lambda_{2}$ $=\frac{-\left(\bar{\beta}\bar{\alpha}a_{\mathrm{max}}^{\text{SE}}-\beta-\alpha\right)-\sqrt{\left(\bar{\beta}\bar{\alpha}a_{\mathrm{max}}^{\text{SE}}-\beta-\alpha\right)^{2}-4\beta\alpha}}{2}=-1$ (67)
$a_{\mathrm{max}}^{\text{SE}}$ $=\frac{1+\beta\alpha+\beta+\alpha}{\bar{\beta}\bar{\alpha}}=\frac{(1+\alpha)(1+\beta)}{\bar{\beta}\bar{\alpha}}$ (68)
For all, $a>a_{\mathrm{max}}^{\text{SE}}$, it follows that $\rho(\bar{A}_{\mathrm{SE}})=|\lambda_{2}|>1$ and the system is unstable.
Corollary 4.5.1.
In the sub-threshold regime (i.e. $\vartheta=\infty$), for $\tau_{u},\tau_{w},\Delta t>0$ and $\alpha=e^{-\frac{\Delta t}{\tau_{u}}}$ and $\beta=e^{-\frac{\Delta t}{\tau_{w}}}$, the Symplectic-Euler discretized adLIF neuron (SE-adLIF) is asymptotically stable for all $a\in~{}(-1,a_{\mathrm{max}}^{\text{SE}})$ with $a_{\mathrm{max}}^{\text{SE}}=\frac{(1+\beta)(1+\alpha)}{(1-\beta)(1-\alpha)}$.
This is a consequence of Lemma 4.4 and Lemma 4.5.
4.6 Proof of Theorem 2.1
Theorem 2.1 states that for each choice of $\tau_{u}>0$, $\tau_{w}>0$, and each intrinsic frequency $f\in[0,f_{N}]$, there is a unique parameter value $[a^{\text{SE}}_{1},a^{\text{SE}}_{2}]$ such that the SE-adLIF neuron with these parameters has intrinsic frequency $f$ and vice versa, while being asymptotically stable in the sub-threshold regime. In other words, the neuron model can exhibit the full range of intrinsic frequencies for any setting of $\tau_{u}>0$, $\tau_{w}>0$ in a stable manner. In the following we proof Theorem 2.1.
Proof.
We first show that for an arbitrary intrinsic frequency $f\in(0,f_{N}]$, there exists a value $a$ such that an SE-adLIF neuron with arbitrary $\tau_{u}>0$ and $\tau_{w}>0$ oscillates with $f$. To show this, we consider an SE-adLIF neuron in the sub-threshold regime with an arbitrary choice of $\tau_{u}>0$ and $\tau_{w}>0$. Let $g_{\tau_{u},\tau_{w}}:[a^{\text{SE}}_{1},a^{\text{SE}}_{2}]\rightarrow[0,f_{N}]$ be the function that maps parameter $a$ of the neuron to the neuron’s intrinsic frequency $f$. We prove in the following that $g_{\tau_{u},\tau_{w}}$ is a bijection.
As proven in the Lemma 4.4, in the range $a\in(a_{1}^{\text{SE}},a_{2}^{\text{SE}})$, with $a_{1}^{\text{SE}}=\frac{\left(\sqrt{\beta}-\sqrt{\alpha}\right)^{2}}{\bar{\beta}\bar{\alpha}}$ and $a_{2}^{\text{SE}}=\frac{\left(\sqrt{\beta}+\sqrt{\alpha}\right)^{2}}{\bar{\beta}\bar{\alpha}}$, $\bar{A}_{\mathrm{SE}}$ has complex eigenvalues $\lambda_{1,2}$ given by Eq. (52).
Let $a\in[a_{1}^{\text{SE}},a_{2}^{\text{SE}}]$ for an arbitrary choice of $\tau_{u}>0$ and $\tau_{w}>0$, and $\lambda_{1}(a)$ be the complex eigenvalue of $\bar{A}_{\mathrm{SE}}$ as function of $a$ with positive argument $arg(\lambda_{1})\geq 0$.
We claim there is a natural bijection $\phi(a)$ between $[a_{1}^{\text{SE}},a_{2}^{\text{SE}}]$ and $[0,\pi]$. We first show that $\cos(\phi(a))$ is a bijection between $[a_{1}^{\text{SE}},a_{2}^{\text{SE}}]$ and $[-1,1]$. The bijection to $[0,\pi]$ follows directly from $\arccos(x):[-1,1]\to[0,\pi]$ defined as a bijective function on its principal values $[0,\pi]$.
We have the trigonometric relation $\Re(\lambda_{1}(a))=r\cos(\phi(a))$ with $r=|\lambda_{1}(a)|=\sqrt{\beta\alpha}$. $\cos(\phi(a)))=\frac{\Re(\lambda_{1}(a))}{\sqrt{\beta\alpha}}$ is a bijection between $[a_{1}^{\text{SE}},a_{2}^{\text{SE}}]$ and $[-1,1]$. $\cos(\phi(a)))$ is surjective, since it is continuous in $a$, $\cos\left(\phi(a_{1}^{\text{SE}}))\right)=1$, and $\cos\left(\phi(a_{2}^{\text{SE}}))\right)=-1$. $\cos(\phi(a))$ is injective, as for $a\in(a_{1}^{\text{SE}},a_{2}^{\text{SE}})$, $\Re(\lambda_{1}(a))$ is a strictly decreasing continuous function, which follows from the fact that its derivative $\Re^{\prime}(\lambda_{1}(a))=-\frac{\bar{\beta}\bar{\alpha}}{2}<0$ for all $a\in(a_{1}^{\text{SE}},a_{2}^{\text{SE}})$, and $r=\sqrt{\beta\alpha}$ is a positive constant.
It follows that $\phi(a)=\arccos{\frac{\Re\left(\lambda_{1}(a)\right)}{\sqrt{\beta\alpha}}}$ defines a bijective function between $[a_{1}^{\text{SE}},a_{2}^{\text{SE}}]$ and $[0,\pi]$, with $\phi(a_{1}^{\text{SE}})=\arccos(1)=0$ and $\phi(a_{2}^{\text{SE}})=\arccos(-1)=\pi$.
Since the frequency in Hertz is defined by $f=\frac{\phi}{2\pi\Delta t}$, we have shown that
$g_{\tau_{u},\tau_{w}}:[a^{\text{SE}}_{1},a^{\text{SE}}_{2}]$ $\to[0,f_{N}]$ (69)
$a$ $\mapsto\frac{\phi(a)}{2\pi\Delta t}$
is a bijective function with $g_{\tau_{u},\tau_{w}}(a_{1}^{\text{SE}})=0$ and $g_{\tau_{u},\tau_{w}}(a_{2}^{\text{SE}})=f_{N}$, the Nyquist frequency.
Hence, we have shown that for an arbitrary intrinsic frequency $f\in(0,f_{N}]$, there exists a value $a$ such that an SE-adLIF neuron with arbitrary $\tau_{u}>0$ and $\tau_{w}>0$ oscillates with $f$.
Second, we have to show that an SE-adLIF neuron with arbitrary parameters $(a,\tau_{u}>0,\tau_{w}>0)$ and intrinsic oscillation frequency $f\in(0,f_{N}]$ is asymptotically stable with decay rate $r=\sqrt{\alpha\beta}<1$ where $\alpha=e^{-\frac{\Delta t}{\tau_{u}}}$ and $\beta=e^{-\frac{\Delta t}{\tau_{w}}}$. Above we have shown that for parameter $a\in[a_{1}^{\text{SE}},a_{2}^{\text{SE}}]$ a bijective mapping to each frequency $f\in(0,f_{N}]$ exists. The asymptotic stability follows from this fact in combination with Lemma 4.4.
4.7 Stable ranges for intrinsic frequencies of EF-adLIF
We consider the parameterizations of EF-adLIF and SE-adLIF neurons where the discretized transition matrix $\bar{A}_{\mathrm{SE}}$ has complex eigenvalues. In that case, the neuron exhibits oscillations of intrinsic frequency $f$ determined by the angle $\phi=\text{arg}\left(\lambda_{1}\right)$ of the complex eigenvalue $\lambda_{1}$, see also Sections 4.4 and 4.5. This angle determines the angle of ’rotation’ of the state for each time step, from which we can infer the intrinsic frequency by $f=\frac{\phi}{2\pi}f_{S}$ where $f_{S}=\frac{1}{\Delta t}$ is the sampling frequency. We define the Nyquist frequency as half the sampling frequency $f_{N}=\frac{f_{S}}{2}$.
Lemma 4.6.
EF-adLIF neurons can oscillate with intrinsic frequencies $f$ bounded by $f\in[0,\frac{f_{N}}{2})$
Proof.
We proved in Lemma 4.2 that in the range $a\in(a_{0}^{\text{EF}},a_{\max}^{\text{EF}})$, with $a_{0}^{\text{EF}}=\frac{(\alpha-\beta)^{2}}{4\bar{\alpha}\bar{\beta}}$ and $a_{\max}^{\text{EF}}=\frac{1-\alpha\beta}{\bar{\alpha}\bar{\beta}}$, the EF-adLIF neuron is asymptotically stable and has complex eigenvalues $\lambda_{1,2}$ given by Eq. (30).
$\lambda_{1}$ is restricted to the right half-plane of the complex plane, which follows directly from the fact that, by definition $\Im(\lambda_{1})>0$ and $\Re(\lambda_{1})=\frac{\beta+\alpha}{2}>0$, for all $\tau_{u}>0$ and $\tau_{w}>0$. Hence, the argument is restricted to $0\leq\phi<\frac{\pi}{2}$. Since the frequency in Hertz is defined by $f=\frac{\phi}{2\pi\Delta t}$, this results in an upper bound on the oscillation frequency of $f<\frac{f_{N}}{2}$ with $f_{N}$ the Nyquist frequency.
Note, that this upper bound is approached in the limit $\tau_{u},\tau_{w}\rightarrow 0$, but the maximum frequency is much lower for realistic values of $\tau_{u}$ and $\tau_{w}$, as shown in Fig. 3d. ∎
Lemma 4.7.
For EF-adLIF, as $\tau_{u}$ and $\tau_{w}$ increase, the stable frequency bandwidth of EF-adLIF asymptotically converges towards 00.
Proof.
Recall that $\alpha=e^{-\frac{\Delta t}{\tau_{u}}}$ and $\beta=e^{-\frac{\Delta t}{\tau_{w}}}$ for $\tau_{u}>0$ and $\tau_{w}>0$. We evaluate $\Im(\lambda_{1})$ at the stability boundary $a_{\max}^{\text{EF}}$ and obtain $\Im\left(\lambda_{1}(a_{\max}^{\text{EF}})\right)=\frac{\sqrt{-(\alpha+\beta)^{2}+4}}{2}$. At this point the modulus is $r=1$ and the maximum radial frequency is thus,
$\phi_{\max}=\arcsin{\frac{\sqrt{-(\alpha+\beta)^{2}+4}}{2}}.$ (70)
As $\lim\limits_{\tau\to\infty}e^{-\frac{\Delta t}{\tau}}=1$, it is clear that $\lim\limits_{\tau_{u},\tau_{w}\to\infty}\phi_{\max}=0$. ∎
4.8 Benchmark Datasets and Preprocessing for Section ’Improved performance of SE-discretized adaptive RSNNs’
[…]
4.9 Training and Hyperparameter Search Details for all Tasks
Optimizer and surrogate gradient
We trained the SNNs using the SLAYER surrogate gradient slayer defined by $\frac{\partial S}{\partial v}=\frac{c\alpha}{2\exp\left(\alpha|v|\right)}$ with $\alpha$ and $c$ according to Table 4 for all experiments. We found that careful choice of the scale parameter $c$ is crucial to achieve good performance for all networks in which the SLAYER gradient is used. A too high $c$ results in an exploding gradient, whereas a too small $c$ can result in vanishing gradients. We trained all networks with back-propagation through time with minibatches using PyTorch anselPyTorchFasterMachine2024. We used the ADAM kingma2017adam optimization algorithm for all experiments with $\beta_{1}=0.9$, $\beta_{2}=0.999$ and $\epsilon=10^{-8}$. For the LIF and adLIF models we detached the spike from the gradient during the reset, such that $u[k]=\hat{u}[k]\cdot(1-\text{sg}\left(S[k]\right))$ where sg is the stop-gradient function with $\text{sg}(x)=x$ and $\frac{\partial}{\partial x}\text{sg}(x)=0$.
Loss functions
In all tasks, the last layer of the network consisted of leaky integrator neurons that match the number of classes of the corresponding task, or the number of masses in case of the oscillatory dynamical system trajectory prediction task. For the SHD and BSD tasks, the loss function was given by $L=\text{CrossEntropy}\left(\sum_{t}\text{softmax}(\bm{y}_{t}),\bm{c}\right)$ for one-hot encoded target class $\bm{c}$ and network output $\bm{y}^{t}$ at time step $t$. We discarded the network output of the first $10$ time steps for the calculation of the loss for SHD, whereas for the BSD task we discarded the output of the first $80\%$ of time steps. For SSC, we used the loss function $L=\text{CrossEntropy}\left(\text{softmax}(\sum_{t}\bm{y}_{t}),\bm{c}\right)$. Again, we discarded the network output of the first $10$ time steps. For the ECG dataset, the loss was computed on a per-time-step level as $L=\sum_{t}\text{CrossEntropy}\left(\text{softmax}(\bm{y}_{t}),\bm{c}_{t}\right)$ where $\bm{c}_{t}$ is the label per time step $t$. For the trajectory prediction task, we used the mean-squared-error (MSE) loss over the temporal sequence, $L=\frac{1}{n\times T}\sum^{T}_{t=1}\sum^{n}_{j=1}\left(\overset{*}{y}^{t}_{j}-y_{j}^{t}\right)^{2}$ where $T$ is the number of time steps in the sequence, $n$ the number of masses and $\overset{\ast}{\bm{y}}^{t}$ is the ground truth of the masses’ displacements.
Hyperparameter tuning
The hyperparameters for the adLIF network were tuned using a mixture of manual tuning and the Hyperband algorithm hyperband, mainly on the SHD classification task. Since the search space of hyperparameters of the adLIF model is quite large, we neither ran exhaustive searches on the ranges for time constants $\tau_{u}$ and $\tau_{w}$, nor the ranges of parameters $a$ and $b$ (for details on how we train these parameters see next section). We selected the ranges based on our stability analyses to ensure stable neurons, and on previous empirical results on similar models bittar2022surrogate. For SSC and ECG we manually tuned only the number of neurons, the learning rate and the SLAYER gradient scale $c$ and kept the values for other hyperparameters same as for the SHD task. We found, that the out-of-the-box performance for these hyperparameters was already very competitive, providing a solid choice as a starting point. For the trajectory prediction task, we performed the hyperparameter search using the Hyperband algorithm hyperband. Only the LIF model was highly sensitive to its hyperparameters – we found that the time constants $\tau_{u}$ and the hyperparameters of the SLAYER gradient function were critical for learning. For the LSTM network, we found that the learning rate should be lower than for LIF and adLIF, and similar to jozefowicz2015empirical, we found that for the model to converge, the forget gate of LSTM should be initially biased to one. For EF-adLIF and SE-adLIF, the performances were mostly invariant to hyperparameter changes, so we conserved the hyperparameters found for SHD and only restricted the parameters $a$ to correspond to the frequency bandwidths described in AdLIF networks accurately predict trajectories of dynamical systems.
Reparameterization and initialization
For all models, we trained the time constants $\tau_{u}$ and $\tau_{w}$ (except for LIF, where no $\tau_{w}$ occurs) in addition to the synaptic weights. We reparameterized the time constants via
$\tau_{x}=\tau_{x}^{\text{min}}+\theta_{x}\left(\tau_{x}^{\text{max}}-\tau_{x}^{\text{min}}\right),$ (71)
with $x\in\{u,w\}$, trained parameters $\theta_{x}$ clipped to the interval $\theta_{x}\in[0,1]$ during training and $\tau_{x}^{\text{min}}$ and $\tau_{x}^{\text{max}}$ as hyperparameters according to Table 4. We found this reparameterization useful if the ADAM optimizer is used, since its dynamic adjustment of the learning rates expects all parameters to be roughly in the same order of magnitude, which is usually not the case for joint training of neuron time constants (order of $10^{1}$ to $10^{2}$) and synaptic weights (order of $10^{0}$). For parameters $a$ and $b$ of the adLIF model we applied a similar reparameterization, where
$a$ $=q\hat{a}$ (72)
$b$ $=q\hat{b},$ (73)
with hyperparameter $q$ and trained parameters $\hat{a}$ and $\hat{b}$, clipped to $\hat{a}\in[0,1]$ and $\hat{b}\in[0,2]$ during training. We restrict the range of parameters to $a,b>0$ to avoid instabilities caused by a positive feed-back loop between adaptation variable $w$ and membrane potential $u$. This constraint has also been discussed in deckers2024co. We initialized the feed-forward weights in all models uniformly in the interval $[-\sqrt{\frac{1}{\text{fan}_{\text{in}}}},\sqrt{\frac{1}{\text{fan}_{\text{in}}}}]$ with $\text{fan}_{\text{in}}$ as the number of inbound synaptic feed-forward connections, and all recurrent weights according to the orthogonal method described in saxe2014a with a gain factor of $1$. $\hat{a}$, $\hat{b}$, $\theta_{u}$ and $\theta_{w}$ are initialized uniformly over their respective range, as stated above. Note, that these are neuron-level parameters, i.e. each neuron has individual values of these parameters.
4.10 Details to the dynamical system trajectory prediction task from Fig. 4
As illustrated in Fig. 4a, we consider a system of $n$ masses connected with $n+1$ springs where each mass (except the two outermost) is connected to two other masses by a spring. The rightmost and leftmost masses are connected to one mass and the fixed support each (see schematic in Fig. 4a). The temporal evolution of the displacements of the masses $\bm{x}(t)\in\mathbb{R}^{n}$ can be written using the equation of motion
$M\ddot{\bm{x}}(t)+S\bm{x}(t)=0,$ (74)
where $M\in\mathbb{R}^{n\times n}$ is a diagonal matrix with diagonal entries corresponding to the masses in kg, while $S\in\mathbb{R}^{n\times n}$ corresponds to the matrix of interaction between the masses determined by the spring coefficients. $S$ is given by
$S_{ij}=\begin{cases}s_{i}+s_{i+1}&\text{if }i=j\\ -s_{i}&\text{if }j=i-1\\ -s_{i+1}&\text{if }j=i+1\\ 0&\text{otherwise}\end{cases}$ (75)
[…]
$\underbrace{\begin{pmatrix}I&\bm{0}\\ \bm{0}&M\end{pmatrix}}_{A}\begin{pmatrix}\dot{\bm{x}}(t)\\ \dot{\bm{v}}(t)\end{pmatrix}=\underbrace{\begin{pmatrix}\bm{0}&I\\ -S&\bm{0}\end{pmatrix}}_{B}\begin{pmatrix}\bm{x}(t)\\ \bm{v}(t)\end{pmatrix},$ (76)
with $\bm{0}$ and $I\in R^{n\times n}$ representing the zero-valued and the identity matrices respectively. The system has a homogeneous solution of the form
$\begin{pmatrix}\bm{x}(t)\\ \bm{v}(t)\end{pmatrix}=\exp{(A^{-1}Bt)}\begin{pmatrix}\bm{x}(0)\\ \bm{v}(0)\end{pmatrix},$ (77)
where $\bm{x}(0),\bm{v}(0)\in\mathbb{R}^{n}$ correspond to the initial conditions of displacements and velocities of the masses respectively.
[…]
The chosen range of spring coefficients resulted in eigenfrequencies of the systems between approximately $2$ to $32$Hz.
For Fig. 4g, we trained our models on different systems with increasing minimal and maximal frequency. Since the frequency of oscillations is determined by $S$ and the spring coefficients, we seek to associate a range of frequencies with a range of spring coefficients from which to sample system parameters. The theoretical frequency bandwidth associated with Eq. (74) can be approximated by calculating the eigenfrequency of the system when all springs are set to the same coefficient $s$, resulting in the matrix $\hat{S}$ of spring coefficients
$\hat{S}_{ij}=\begin{cases}2s&\text{if }i=j\\ -s&\text{if }j=i-1\\ -s&\text{if }j=i+1\\ 0&\text{otherwise}.\end{cases}$ (78)
In the case of unit masses, we can formulate an eigenvalue problem from the system defined by Eq. (74) as
$\hat{S}\bm{v}_{j}=\omega^{2}_{j}\bm{v}_{j},$ (79)
where $\bm{v}_{j}$ is the eigenvector associated to the eigenvalue $\omega^{2}_{j}$ and $\omega_{j}$ is the corresponding radial frequency. $\hat{S}$ is a tridiagonal Toeplitz matrix, as such the $j$-th eigenvalue associated to mass $j$ has a closed form solution kulkarni1999eigenvalues
$\omega^{2}_{j}=4s\sin^{2}{\left(\frac{j\pi}{2\left(n+1\right)}\right)},$ (80)
where $s$ is the spring coefficient and $n$ the number of masses. The maximal eigenvalue is thus given by
$\omega^{2}_{\max}=4s\sin^{2}{\left(\frac{n\pi}{2(n+1)}\right)}\approx 4s,$ (81)
and the radial frequency by
$\omega_{\max}\approx 2\sqrt{s}.$ (82)
The minimal eigenvalue corresponds to
$\omega^{2}_{\min}=4s\sin^{2}{\left(\frac{\pi}{2(n+1)}\right)}\approx 4s\left(\frac{\pi}{2(n+1)}\right)^{2},$ (83)
and the radial frequency is thus given by
$\omega_{\min}\approx\frac{\pi}{n+1}\sqrt{s}.$ (84)
The range of spring coefficients can be determined by setting $\omega_{\min}$ (resp. $\omega_{\max}$) to the desired minimum (resp. maximum) radial frequency and solving for $s$.
[…]
4.11 Details to simulations in Fig. 5
[…]
4.12 Details to the Burst Sequence Detection (BSD) task
[…]
In an input sample $X\in[0,1]^{T\times N}$, $x_{t,n}=1$ indicates a spike of input neuron $n$ at time $t$ ($x_{t,n}=0$ if there is no spike). We generated the spike trains $X$ for input neurons for a given class $c$ as follows. Each $x_{t,n}$ was drawn from a Bernoulli distribution with $p(x_{t,n}=1|c)$ that was obtained as follows. Bursts were modeled as brief, smooth increases in spike probability at the corresponding times $t^{n}$. We achieved this by a Gaussian function $f(t,t^{n})=\exp\left(\frac{-(t-t^{n})^{2}}{4}\right)$, yielding high spike probability for time steps close to $t^{n}$, and lower spike probability further away. From this function, we computed the final spike probability $p(x_{t,n}=1|c)=\frac{f(t,t^{n})}{\max_{k\in[0,T]}f(k,t^{n})}\cdot 0.75+0.05$, interpolating between the minimum and maximum spike probabilities of $0.05$ ($50$ Hz) and $0.8$ ($800$ Hz) respectively. This way we ensured that the spike probability went up to $0.8$ during a burst and was approximately $0.05$ otherwise to mimic background noise. This procedure was repeated for each individual data sample $X$ to obtain the dataset. We randomly chose a class for each new sample, such that the class cardinalities in the dataset were roughly but not exactly balanced. The whole data set consisted of $8000$ samples. We held out $10\%$ ($800$ samples) of this dataset for validation and $20\%$ ($1600$ samples) for testing. For the experiment with different numbers of classes (Fig.6b), we only increased the number of classes, but did not increase the number of samples in the dataset. This resulted in increased difficulty for larger numbers of classes, since less samples per class were present in the training data.
4.13 Optimization-based feature visualization
The visualization of important input features for the trained network shown in Fig. 6 was obtained using optimization-based feature visualization as defined in erhan2009visualizing. We applied the same sampling method to four different networks, one adLIF and one LIF network trained on BSD and one adLIF and one LIF trained on SHD. For both cases we used the same algorithm, but with different parameters, which are listed in Table 2. For both tasks, we selected network instances that achieved high accuracy. The sampling procedure was performed as follows. First, we initialized a random input example $X^{0}\in\mathbb{R}^{T\times C}$ with $C$ as the input dimension and $T$ the input length (number of time steps), both of the same dimension as the data the network was trained on. Here, each $x^{0}_{t,c}$ was drawn from the uniform distribution $\mathcal{U}(0,1)$. At each iteration $k$, the sample was passed through the network and the loss was calculated akin to the loss function used during training on the specific dataset (SHD or BSD) to obtain the gradient $G_{k}=\nabla_{X}L(X,c)|_{X^{k}}$, with respect to a pre-defined target class $c$. This gradient was normalized to obtain $\Delta X^{k}=\eta\frac{G_{k}}{\zeta_{k}}$ with normalization factor $\zeta_{k}=\max\left\{|G_{k}|_{\text{max}},\epsilon\right\}$ using an $\epsilon=10^{-6}$ for numerical stability, and a step size $\eta$. $|\cdot|_{\text{max}}$ denotes the maximum norm returning the highest absolute value. After the gradient update, we clipped the sample $X^{k+1}$ to the positive range, as the data from both datasets, SHD and BSD, is all-positive. We then applied Gaussian smoothing on the data sample $X^{k+1}$. In the BSD case we applied a $1$D smoothing along the time axis using a Gaussian kernel $g(s)=\frac{1}{\sqrt{2\pi}\sigma}\exp\left(-\frac{s^{2}}{2\sigma^{2}}\right)$ via $X^{k+1}\leftarrow\nu X^{k+1}+\gamma(g*X^{k+1})$. Decay $\nu\in[0,1]$ and smoothing coefficient $\gamma\in[0,1]$ are hyperparameters, * denotes a convolution operation. The $\sigma$ parameter of the Gaussian kernel was linearly decayed from $\sigma_{\text{init}}$ to 00 over the $n_{\text{iter}}$ iterations to decrease the effective regularization applied by the kernel smoothing. For the SHD case we applied a $2$D smoothing to the data sample over both the time dimension and the spatial dimension. For the BSD case, we normalized the sample $X^{k+1}$ with $X^{k+1}\leftarrow X^{k+1}\frac{\mu_{\text{all}}}{\mu_{k+1}}$, where scalar $\mu_{\text{all}}$ is the mean spike rate over all data samples and input neurons in the data, and $\mu_{k+1}$ the mean of sample $X^{k+1}$ over both the time and the input neuron axes. For the SHD setup, we ignored the loss for the first $100$ sequence time steps (compare with grey shaded area in Fig. 6e, whereas for the BSD case we ignored the loss for the first $80\%$ of time steps, which was also done during training of the network on the dataset (see Section 4.9).
BSD SHD
$n_{\text{iter}}$ $400$ $200$
$\eta$ $0.1$ $0.1$
$\nu$ $0.891$ $0.891$
$\gamma$ $0.1$ $0.1$
$\sigma_{\text{init}}$ $5$ $5$
Table 2: Hyperparameters for optimization-based feature visualization
4.14 Hyperparameters
We present hyperparameters that remained unchanged for all tasks and models in Table 3 and the task- and model-specific hyperparameters in Table 4.
$\vartheta$ $1$
$\Delta t$ $1$
FF Init. $\mathcal{U}\left(-\sqrt{\frac{1}{\text{fan}_{\text{in}}}},\sqrt{\frac{1}{\text{fan}_{\text{in}}}}\right)$
Rec Init. orthogonal saxe2014a
Gradient clip val. $1.5$
Table 3: Global hyperparameters Gradient clip val.: Gradient norm to which the gradient was rescaled if exceeded. FF Init: initialization of feed-forward synaptic weights. Rec Init: Initialization of recurrent synaptic weights.
lr # neurons # layers q $\alpha$ (SLAYER) c (SLAYER) Ep. [$\tau_{u}^{\text{min}},\tau_{u}^{\text{max}}$] [$\tau_{w}^{\text{min}},\tau_{w}^{\text{max}}$] dropout $\tau_{\text{out}}$ batch size SHD SE $0.01$ $128/360$ $1/2$ $120$ $5$ $0.4$ $300$ $[5,25]$ $[60,300]$ $15\%$ $15$ $256$ EF $0.01$ $360$ $2$ $60$ $5$ $0.4$ $300$ $[5,25]$ $[60,300]$ $15\%$ $15$ $256$ LIF $0.01$ $360$ $2$ $5$ $0.1$ $300$ $[5,150]$ $15\%$ $15$ $256$ SSC SE $0.006$ $720$ $2$ $120$ $5$ $0.4$ $40$ $[5,25]$ $[60,300]$ $15\%$ $15$ $256$ EF $0.006$ $720$ $2$ $60$ $5$ $0.4$ $40$ $[5,25]$ $[60,300]$ $15\%$ $15$ $256$ LIF $0.006$ $720$ $2$ $5$ $0.1$ $40$ $[5,150]$ $15\%$ $15$ $256$ ECG SE $0.01$ $36$ $1/2$ $120$ $5$ $0.2$ $400$ $[5,25]$ $[60,300]$ $15\%$ $3$ $64$ EF $0.01$ $36$ $1$ $60$ $5$ $0.2$ $400$ $[5,25]$ $[60,300]$ $15\%$ $3$ $64$ LIF $0.01$ $36$ $1$ $5$ $0.1$ $400$ $[5,150]$ $15\%$ $3$ $64$ BSD SE $0.01$ $512$ $1$ $120$ $5$ $0.4$ $400$ $[5,25]$ $[60,300]$ $0\%$ $15$ $128$ LIF $0.006$ $510$ $1$ $5$ $0.2$ $400$ $[5,50]$ $0\%$ $15$ $128$ spring-mass SE/EF $0.01$ $[25,3200]$ $1$ $228$/$65$ $5$ $0.4$ $200$ $[5,25]$ $[60,300]$ $0\%$ $[1,20]$ $256$ LIF $0.01$ $[27,3202]$ $1$ $10$ $0.5$ $200$ $[1,25]$ $0\%$ $[1,20]$ $256$ LSTM $0.001$ $[13,1600]$ $1$ $200$ $0\%$ $[1,20]$ $256$
Table 4: Task- and model-specific hyperparameters lr: learning rate for ADAM optimizer, Ep.: number of training epochs, $\tau_{\text{out}}$: Membrane time constants of output layer (leaky integrator), dropout: dropout rate
5 Data Availability
The SHD and SSC datasets are publicly available under https://zenkelab.org/resources/spiking-heidelberg-datasets-shd/. The QTDB ECG dataset is publicly available under https://physionet.org/content/qtdb/1.0.0/. However, we used the preprocessed files from yinAccurateEfficientTimedomain2021.
6 Code Availability
The code is available under the CC BY-SA 4.0 license at the repository https://github.com/IGITUGraz/SE-adlif.
7 Acknowledgements
This work has been supported by the “University SAL Labs” initiative of Silicon Austria Labs (SAL) and its Austrian partner universities for applied fundamental research for electronic based systems, by the European Community’s Horizon 2020 FET-Open Programme, grant number 899265, ADOPD, and by NSF EFRI grant #2318152. We thank Felix Effenberger and Sebastian Otte for valuable feedback on the manuscript.
Declarations
The authors declare no conflict of interest.
Author Contribution
M.B., R.F., and R.L. conceived the idea and proposed the research, M.B., R.F. and S.S. carried out experiments, M.B. and R.F. performed theoretical analyses, M.B., R.F., and R.L. wrote the paper.
[…]
Appendix A Stability issues of EF-adLIF neurons
In all experiments with SE-adLIF neurons, we used the same parameter range as in Fig. 3a: $a\in[0,120]$, $\tau_{u}\in[5,25]$ ms and $\tau_{w}\in[60,300]$ ms. As shown in the same panel, this parameter range is unstable for EF-adLIF, hence we reduced the range of $a$ to $[0,60]$ for all experiments with EF-adLIF to ensure a decay rate $r<1$ for all parameter configurations. If the larger range of $a\in[0,120]$ is used for EF-adLIF, we can observe instabilities during training, resulting in significant degradation of performance. We report the performance drop on the ECG dataset in Table 5. Within these experiments, we observed exploding gradients during training in many of the runs with the extended range of $a$, which were not present for SE-adLIF with the same parameter range. Partially, these exploding gradients can be alleviated by using gradient clipping, resulting in slightly better performance. This experiment demonstrates the sensitivity of EF-adLIF stability to choices of parameter ranges. Note, that this sensitivity can directly be linked to the relation between neuron parameters and eigenvalues of $\bar{A}_{\mathrm{EF}}$, as shown in Fig. 3.
range of $a$ Grad. clip. Test Acc. [%]
$[0,60]$ $87.10\pm 0.46\%$
$[0,120]$ $70.66\pm 24.90\%$
$[0,120]$ $55.62\pm 24.84\%$
Table 5: Instabilities of EF-adLIF for the ECG task. Test accuracy for different ranges of parameter $a$ for EF-adLIF. Grad. clip. denotes rescaling the gradient to a norm of $1.5$ if exceeded. Table shows mean and std. dev. over $5$ runs. Chance level in this task is $1/6=16.67\%$. Note the high variance of accuracies in the less constrained cases.
Appendix B Neural heterogeneity of trained adLIF SNNs
In gerstner2014neuronal, the authors show that adLIF models can, depending on their parametrization, account for many experimentally observed neocortical neuron spiking behaviors. It remained an open question whether these different neuronal dynamics and firing patterns could be beneficial in sequence processing tasks. To that end, we analyzed adLIF networks trained on the SHD task and observed a vast heterogeneity in the evolved neuron parameterizations, see Fig. 8.
Figure 8: Heterogeneity of adLIF neurons in a network trained on SHD For the analysed network we used the same architecture as for the results reported in Table 1. The network achieved an accuracy of $97.08\%$ on the SHD data set. a) Comparison of the amount of underdamped and overdamped neurons in the network after training. b) Scatter plot of underdamped neurons with respect to their frequency $f$ (x-axis), decay rate $r$ (y-axis) and spike-triggered adaptation parameter $b$ (color). Histogram shows the marginal distribution, summed over the y-axis. c) Scatter plot of all neurons with respect to their intrinsic frequency $f$ and parameter $b$. Histograms show marginal distributions over the corresponding axes. d) Density plot of the distribution of effective time constants $\tau_{\mathrm{eff}}=-\frac{\Delta t}{\ln r}$ for underdamped (left) and overdamped (right) neurons. Horizontal lines denote the mean. e) Responses of different classes of adLIF neuron parameterizations to different input stimuli: pulse, step and ramp. Parameter $b$ controls the strength of spike-triggered adaptation. The four classes are OL: overdamped with low $b$, equivalent to a LIF neuron, OH: overdamped with high $b$, UL: underdamped with low $b$, and UH: underdamped with high $b$. The plots show the membrane potential over time in response to the input current in the first row. Vertical lines show output spikes of the neuron. All neuron parameterizations were taken from the same network as panels a-d.
We categorized the observed neurons based on features derived from their parameters and properties. One first distinction that can be made is whether the neurons exhibit integratory (overdamped, type I) or oscillatory membrane potentials (underdamped, type II), shown in Fig. 8a. We found that most of the neurons ($87\%$) showed underdamped behavior. Most of these underdamped neurons converged to rather slow frequencies of $f<50$ Hz, Fig. 8b. As shown in Fig. 2d and e, these slow-frequency neurons can act as feature detectors for input variations much slower than the spike rates. Another categorization can be made based on the parameter $b$: independent of oscillatory properties, this parameter determines the strength of feed-back from neuron spikes to the adaptation variable $w$. Interestingly, neurons with diverse degrees of such spike-triggered adaptation occured across the entire frequency range (color in Fig. 8b). The distribution of $b$ was rather uniform, with two peaks where we clipped the parameter at the minimum and maximum range, representing weak and strong spike-triggered adaptation respectively. Interestingly, overdamped neurons acted on much longer timescales (see Fig. 8d) than underdamped ones. This figure shows the distribution of the effective time constant $\tau_{\mathrm{eff}}=-\frac{\Delta t}{\ln r}$, which is analogous to the membrane time constant of LIF neurons.
The heterogeneous combinations of the spike-triggered adaptation governed by parameter $b$ and the oscillatory behavior governed by parameter $a$ result in diverse non-linear super- and sub-threshold neuron dynamics (Fig. 8e). Based on the overdamped-underdamped split and the distinction into high and low $b$-values, we selected four neurons from the network for further analysis. These four classes are hence given by overdamped-low-$b$ (OL) neurons, that are equivalent to LIF neurons, overdamped-high-$b$ (OH), underdamped-low-$b$ (UL) and underdamped-high-$b$ (UH) neurons. For example, overdamped neurons with high $b$ exhibit a hyper-polarization, similar to a relative refractory period after emitting a spike (2nd row, 1st column). A second example of an interesting learned behavior is given by underdamped neurons with low spike-triggered adaptation (class UL), which respond with an immediate short burst of action potentials to a positive step in the input current, but with a delayed short burst to a negative step (3rd row, 2nd column). Another intriguing property is the interference of spike-triggered adaptation and oscillations, in which an output spike of an UH-neuron could cancel the membrane potential oscillation (4th row, 3rd column). Such properties don’t exist in LIF neurons (class OL), which only spike in response to sufficient input current, without any complex dynamics. In summary, our analysis revealed that recurrent networks of adLIF neurons exhibit a diverse set of neuronal dynamics when trained for optimal performance on a temporal classification task. This diversity is reminiscent of the diversity of neuronal behaviors found in biological neuronal networks izhikevich2001resonate; gerstner2014neuronal.
Appendix C Additional experiments on spring-mass system
Figure 9: Ablation study for AdLIF networks accurately predict trajectories of dynamical systems. Comparison of LIF and adLIF networks trained with and without recurrent connections. Degradation time denotes the time when the model MSE reaches the MSE of the constant model.
Appendix D Performance and visualized features of networks of neurons with threshold adaptation
We evaluated recurrent spiking networks with a modified variant of the ALIF neuron yinAccurateEfficientTimedomain2021 on the BSD task. For the case of $20$ classes this model achieved a test accuracy of $86.53\pm 1.01\%$ over $10$ runs. For these experiments, we used the same model as in the paper from Yin et al. yinAccurateEfficientTimedomain2021 except for the following modifications: Neuron time constants $\tau_{m}$ and $\tau_{\text{adp}}$ were trained with the reparametrization described in Section Training and Hyperparameter Search Details for all Tasks in Methods and the initialization was the same as for our other experiments (see Table 4). We applied the optimization-based feature visualization from Section Networks of adLIF neurons tune to high-fidelity temporal features to an ALIF network trained on the $20$-class BSD task and show the resulting feature maps in Fig. 10. The hyperparameters for the ALIF model for the BSD task are shown in Table 6.
Figure 10: Gradient-based feature visualization applied to ALIF models. The feature maps of the ALIF network with highest accuracy on the $20$-class BSD task for a target class of $2$ (left) and $17$ (right).
lr # neurons # layers q $\alpha$ (SLAYER) c (SLAYER) Ep. [$\tau_{u}^{\text{min}},\tau_{u}^{\text{max}}$] [$\tau_{\text{adp}}^{\text{min}},\tau_{\text{adp}}^{\text{max}}$] dropout $\tau_{\text{out}}$ batch size $0.008$ $511$ $1$ $120$ $5$ $0.4$ $400$ $[5,100]$ $[60,200]$ $0\%$ $15$ $128$
Table 6: Hyperparameters of the ALIF model for the BSD task. lr: learning rate for ADAM optimizer, Ep.: number of training epochs, $\tau_{\text{out}}$: Membrane time constants of output layer (leaky integrator), dropout: dropout rate